### v1

# An Evaluation of N-Gram Selection Strategies for Regular Expression Indexing in Contemporary Text Analysis Tasks [Experiment, Analysis & Benchmark]

Abstract.
Efficient evaluation of regular expressions (regex, for short) is crucial for text analysis, and n-gram indexes are fundamental to achieving fast regex evaluation performance. However, these indexes face scalability challenges because of the exponential number of possible n-grams that must be indexed. Many existing selection strategies, developed decades ago, have not been rigorously evaluated on contemporary large-scale workloads and lack comprehensive performance comparisons. Therefore, a unified and comprehensive evaluation framework is necessary to compare these methods under the same experimental settings. This paper presents the first systematic evaluation of three representative n-gram selection strategies across five workloads, including real-time production logs and genomic sequence analysis. We examine their trade-offs in terms of index construction time, storage overhead, false positive rates, and end-to-end query performance. Through empirical results, this study provides a modern perspective on existing n-gram based regular expression evaluation methods, extensive observations, valuable discoveries, and an adaptable testing framework to guide future research in this domain. We make our implementations of these methods and our test framework available as open-source at https://github.com/mush-zhang/RegexIndexComparison.
1. Introduction
Regular expressions are a foundational tool for text pattern matching, powering critical applications such as real-time log analysis (Rouillard, 2004), genomic sequence alignment (Arslan, 2005), and web information retrieval (De Bra and Post, 1994). However, as datasets grow in their size, the computational cost of brute-force matching becomes prohibitive. To address this issue, n-gram indexing has been widely adopted to accelerate regex processing by pre-filtering candidate text regions using selected n-grams (Ogawa and Matsuda, 1998; Cox, 2012). Despite its widespread adoption, the scalability of this approach hinges on a critical problem: how to select the optimal set of n-grams to index, balancing trade-offs between index size, construction time, and query accuracy.
Existing solutions, such as frequency-based selection (prioritizing common n-grams) (Cho and Rajagopalan, 2002), coverage-optimized methods (minimizing redundant matches) (Hore et al., 2004), and heuristic-driven strategies like LPMS (Tsang and Chawla, 2011), were proposed decades ago. For instance, the FREE (Cho and Rajagopalan, 2002) and BEST (Hore et al., 2004) methods, introduced in the early 2000s, focus on minimizing computational space overhead and index size for memory-constrained systems. Despite being an active area of research, surprisingly, the performance of these methods on modern hardware is lacking for three reasons. The absence of a comprehensive, unbiased, and systematic understanding of these methods is due to three main reasons. First, existing methods have not been evaluated within the same experimental settings (such as the same hardware or using the same B+-Tree implementation), leading to incomplete comparisons. Second, prior work has only compared specific metrics and on a limited number of datasets. Third, the absence of a standardized and inclusive testing framework hinders the investigation and testing of methods in practical implementations. Consequently, the lack of a modern comparison of these past methods has led to the adoption of these legacy strategies in recent studies (Qiu et al., 2022; Chockchowwat et al., 2022) without re-evaluating their assumptions. These older methods may not be appropriate when dealing with larger datasets (now common in applications domains such as genenomics, IoT/sensor, and log analytics) as it stresses their n-gram selection methods. For example, as we will see later, BEST’s quadratic runtime complexity becomes prohibitive on even a small 10GB text corpus, forcing practitioners to downsample data or abandon indexing altogether. In fact, n-gram indexing used by Postgres (Christopher Kings-Lynne and Korotkov, [n.d.]) and GitLab code search (Peterse, 2016) abandon the n-gram selection strategies and build the index with all n-grams for only trigrams (i.e. $n=3$) or for several fixed values $n$ (for instance, indexing all bigrams and trigrams). In summary, there are three key limitations that motivate our work.
  1. M.1
    Lack of Comprehensive Comparison. Prior evaluations focus on specific metrics (e.g., false-positive rate) or evaluate their method using only a single workload, failing to capture the diverse demands of real-world workloads. For example, genomic databases prioritize minimizing false positives due to a large number of repeated queries on the same dataset, while log analysis prioritizes rapid index rebuild overhead for large scale streaming data analysis. No study has systematically compared and contrasted the three key methods FREE, BEST, and LPMS across wide-ranging scenarios.
  2. M.2
    Outdated Resource Assumptions. Existing methods often optimize assuming limited memory, investing significant efforts to reduce memory overhead and index sizes. Additionally, some approaches focus on scenarios where the index and/or dataset cannot fit in memory and use the IO cost as an optimization parameter (Kim et al., 2010). However, in many modern hardware settings, abundant memory is often available, and it is also important to consider this case.
  3. M.3
    Fine-grained empirical analysis. Beyond evaluating the existing methods on a comprehensive set of common metrics, we also conduct detailed resource usage measurements for each method, providing deeper performance related insights.
Our Contributions. We systematically study three state-of-the-art methods, FREEBEST, and LPMS, across various workloads. Specifically, we make the following contributions.
1. Systematic Method Analysis. We formalize and categorize three state-of-the-art n-gram selection strategies, FREE (frequency-based), BEST (coverage-optimized), and LPMS (linear programming approximation), within a unified framework. This enables direct theoretical comparison of their computational complexities. Additionally, we provide a detailed account of their implementation and design decisions, emphasizing both similarities and subtle differences with other methods, addressing (M.1)
2. Methods and Benchmark Framework Implementation. In the absence of original code, we implemented FREE, BEST, and LPMS as described in the respective papers, using modern C++. While BEST’s original design includes parallelism for multi-core CPUs, FREE and LPMS lack scalable implementations. We modernized all three methods by: (1) redesigning FREE with parallelism to leverage modern multi-core CPUs, and (2) implementing LPMS with Gurobi, which inherently supports multi-threading in its LP solver. Additionally, we introduced an early stopping mechanism to all three methods to configure the maximum number of n-grams selected. We also developed an end-to-end benchmark framework that manages data loading, method selection, configuration settings, regex matching, and result reporting. This framework is modular, easily extensible, and publicly available at the link noted in the abstract. All experiments were conducted on a multi-core machine with large memory, addressing (M.2).
3. Broad Workload Benchmark. To address the lack of comprehensive empirical comparisons on workloads with different characteristics, we evaluate the methods across five workloads. This includes using three legacy benchmarks from prior work and two real-world workloads in contemporary regex matching tasks.
4. Empirical Trade-off Analysis and Guidelines. Our experiments quantify performance across various workloads by measuring index construction time, memory footprint, and index precision, among other metrics (addressing (M.3)). We correlate these results with workload characteristics, such as regex complexity and dataset size, to derive actionable guidelines. For instance, FREE achieves a 92% reduction in index build time compared to BEST with only a 1.2% increase in query latency, making it more suitable for streaming log analysis. These findings challenge the necessity of computationally intensive methods for large-scale workloads.
2. Related Work
Efficient indexing of regular expressions is crucial to enhance query performance in large-scale database systems, log analysis, and information extraction applications. Traditionally, regular expressions are evaluated using finite automata, which is computationally expensive due to backtracking and state transitions. Since regex queries operate on string datasets, n-gram-based indexing techniques have been widely adopted to accelerate regex evaluation by pre-filtering candidate matches before full regex matching. N-gram selection techniques are critical for regex indexing, as they impact both query filtering accuracy and index storage overhead.
To maximize regex evaluation performance before the introduction of n-gram indexing techniques or when such indexing is not feasible, early research focused on string literal filtering to optimize regex performance. A suffix-based string matching heuristic was presented in (Boyer and Moore, 1977), laying the foundation for later suffix-tree-based regex indexing methods. Suffix-based indexing was then proposed for efficient pattern matching over large datasets (Gusfield, 1997). Literal filtering, or prefix heuristics for approximate string matching, was introduced in (Wu and Manber, 1992; Clarke, 1996). These studies demonstrated that prefix-based early termination can significantly reduce regex search time. Later works proposed filtering with all discriminative string literals in addition to prefixes and suffixes in the regexes (Wang et al., 2019; Zhang et al., 2023).
N-gram indexing was introduced to further improve regex query performance by precomputing substrings of length n as index keys. Besides being used for pattern matching, n-gram indexing is also common in approximate string searching. Early works compared different n values and decided to index all trigrams. Other works index n-grams with different n values for various workloads. The database community has conducted extensive research on effective indexing strategies for regular expression queries. While existing works primarily focus on improving regex query performance through optimized index structures, relatively fewer studies have explored n-gram key selection strategies for regex indexing. N-gram indexing has also been used for the closely related problem of indexing regular expressions (instead of data) to find out which regexes match a given input string (Chan et al., 2003). Similar ideas have also been used for indexing data to speed up regular path query evaluation (Li et al., 2001).
The first known works discussing n-gram selection for regex indexing is FREE (Cho and Rajagopalan, 2002). The high-level idea of selecting a covering n-gram set was proposed in a study of Asian language indexing (Ogawa and Matsuda, 1998). The concept of a covering set was further developed by Kim et al. (Kim et al., 2010) for a more efficient n-gram selection method considering I/O cost. The analogy between the set-covering problem and the n-gram selection problem was formalized in BEST (Hore et al., 2004), where the authors presented a near-optimal algorithm to select variable-length n-grams considering index size constraints. Later work, LPMS (Tsang and Chawla, 2011), combined the solution formulations of FREE and BEST and reformulated the n-gram selection problem into linear programs. In this work, we compare the n-gram selection and indexing methods of FREE, BEST, and LPMS for in-memory workloads and indices.
3. Problem Definition
In this section, we discuss the n-gram selection problem and provide formal definitions for some common terms.
Definition 3.1.
N-Gram For a given finite alphabet $\Sigma$, an n-gram $g$ is a sequence of $n$ characters $g=g_{1}g_{2}\cdots g_{n}$ where $g_{i}\in\Sigma$.
We use $|g|$ to denote the length of the n-gram $g$. A literal is a string from the set $\Sigma*$, where * is the standard closure operator. A common operation on a regex query $q$ is to identify the maximal literal components from the regex. For a literal $l$, we will use $G(l)$ to denote the set of all possible substrings of $l$. The set of all possible n-grams of a regex $q$ is $G(r)=\bigcup_{\text{maximal }l\in q}G(l)$.
Example 3.2.
Regex queries often times look for patterns as a sequence of characters. The regex <ahref=("|’).*ZZZ\. pdf("|’)> from a workload of regex queries over webpages matches a URL pattern. The set of literal components in the example regex are: {<ahref=, ZZZ.pdf, >}.
A workload $W=(Q,D)$ consists of the set of regex query $Q=\{q_{1},q_{2},\cdots\}$ and dataset $D=\{d_{1},d_{2},\cdots\},d_{i}\in\Sigma*$. The size of the dataset is defined as $|D|=\sum_{d_{i}\in D}|d_{i}|$. Each $d_{i}$ will be referred to as a data record. Next, we define the support of an n-gram.
Definition 3.3.
Support The support $s$ of a literal $g$ in dataset $D$ is the number of elements in $D$ that contains $g$. Similarly, its support in query set $Q$ is the number of individual queries which contains $g$ as a literal.
$s_{D}(g)=\sum_{d\in D}\mathbbm{1}\left[g\in G(d)\right]\quad s_{Q}(g)=\sum_{q\in D}\mathbbm{1}\left[g\in G(q)\right]$
Using the notion of support, we define the selectivity of an n-gram $g$ as follows.
Definition 3.4.
Selectivity The selectivity $c$ of an n-gram $g$ in string set $D$ is the fraction of individual strings in $D$ that contains $g$.
$c_{g}=\frac{s_{D}(g)}{|D|}$
For n-gram selection methods, selectivity is an important parameter when deciding if an n-gram is selected for indexing or not. For the previous example regex that looks for a URL of a PDF file which has a filename end with sub-string ZZZ, by looking for strings with n-grams ZZZ.pdf, intuitively, we can reduce the number of data strings for the exact regex matching. Notice that there are two other literals in the regex, <ahref= and >. Since these two literals almost always occur in every webpage HTML, indexing these two literals might not reduce the number of data records for exact regex matching, that is, high selectivity. Further, it may result in extra computation overhead. Conversely, n-gram that has lower selectivity like ZZZ.pdf can help reduce the overall query time.
However, lower selectivity is not always better, when considering the entire query set, Q. For instance, the n-gram ZZZ.pdf may have low selectivity and effectively filter out irrelevant data points. Since ZZZ is an uncommon character sequence in the English language, it might not appear in other queries. On the other hand, indexing .pdf, which has relatively higher selectivity than ZZZ.pdf, might not reduce as many data points, but it can benefit other queries that look for URLs of PDF files with different names. Thus, there is a trade-off between selecting higher and lower selectivity n-grams.
4. Methods Overview
In this section, we present the overview of the three state-of-the-art n-gram selection methods: FREE, BEST, and LPMS. We will describe their selection strategy and provide a complexity analysis. In Section 4, we summarize and compare the three methods in terms of their source of n-gram, selection criteria in each step, index structure accompanied, and other common configurations.
Table 1. Selection methods summary.
Method N-Gram Source Selection Criteria Selectivity Threshold ($c$) N-Gram Constraint Index Structure Index Size Constraint
FREE (Cho and Rajagopalan, 2002) dataset prefix-free, selectivity 0.1 $2\leq n\leq 10$ Inverted Index -
BEST (Hore et al., 2004) queries & dataset Utility(g) = benefit($g$)/cost($g$) $0.05\leq c\leq 0.1$ B+-Tree Index User defined
LPMS (Tsang and Chawla, 2011) queries & dataset prefix-free and utility optimized - - Inverted Index -
4.1. FREE
4.1.1. Selection Strategy
FREE uses the dataset in the workload as the source of n-gram selection, and selects a prefix-free set of n-grams based on selectivity. Note that FREE does not use the query workload for n-gram selection. The candidate set $G(W)$ of n-grams is $G(W)=\bigcup_{d\in D}G(d)$. For the candidates $G(W)$, FREE decides whether a n-gram will be selected by Usefulness.
Definition 4.1 (FREE).
Usefulness By setting a fixed selectivity threshold $\mathbf{c}$, n-gram $g$ is deemed useful if its selectivity is less than the selectivity threshold, $c_{g}<\mathbf{c}$.
FREE makes two important assumptions when selecting n-grams for indexing.
Assumption FREE-1 0.
FREE assumes that n-grams with high selectivity are less useful.
Assumption FREE-2 0.
With a careful selected selectivity threshold $\mathbf{c}$, FREE assumes that a query without any useful n-grams is rare. Overall workload performance will still improve as most of the regexes can benefit from indexing only useful n-grams.
FREE selects only n-grams that are useful. Since usefulness selection criteria is essentially a selectivity upper bound, if we know that a n-gram $g$ is useful, then any longer n-gram having $g$ as a substring is also useful. For example, if n-gram pdf is useful with a selectivity $c_{g}$, then n-grams such as $g^{\prime}=$pdf and $g^{\prime\prime}=$ pdfg will have selectivity $0\leq c(g^{\prime})<\mathbf{c}$ and $0\leq c(g^{\prime\prime})<\mathbf{c}$, and therefore also useful.
Property FREE: Usefulness 0.
For an n-gram $g=g_{1}g_{2}\cdots g_{n_{1}}$ of size $n_{1}$ having selectivity $c_{g}$, any other longer n-grams of size $n_{2}$
$g^{\prime}=p_{1}\cdots p_{m}g_{1}g_{2}\cdots g_{n_{1}}s_{1}\cdots s_{n_{2}-n_{1}-m}$
with a prefix size $m$ and suffix size $n_{2}-n_{1}-m$, where each character $p_{i},s_{i}\in\Sigma$, we have $0\leq c_{g^{\prime}}\leq c_{g}$, is also useful.
Since the set of all useful n-grams can still be too large, to further reduce the index size, FREE selects only the minimal n-gram among all n-grams with the same prefix. For example, n-grams pdf and pdfg have the same prefixes p, pd, and pdf. If the selectivity of the first two n-grams are not useful while pdf is useful, we index only n-gram pdf, although pdfg is also useful. Thus, the second n-gram selection criterion besides selectivity is minimality.
Definition 4.2 (FREE).
Minimal For an n-gram $g$ in alphabet $\Sigma$ of size $n$ that is useful, it is prefix-minimal among the set of n-grams that has $g$ as a prefix if no prefix substring of $g$ with size smaller than $n$ is useful. An n-gram $g$ is suffix minimal if no suffix substring of $g$ with size smaller than $n$ is useful. N-gram $g$ is prefix-suffix minimal, or pre-suf minimal if no substring of $g$ with size smaller than $n$ is useful.
To effectively reduce the size of n-grams selected, FREE selects the prefix-minimal useful set of n-grams, deriving from the Apriori method of finding the maximal frequent sets in data mining literature (Agrawal et al., 1996). Essentially, it generates all candidate n-grams in increasing size order iteratively. For iteration $i$, it generates the candidate set of n-grams of length $i$ by extending all useless n-grams of length $i-1$ by one character, inserting the useful n-grams in the candidate set into the index, and using the useless ones in the candidate set for iteration $i+1$. This way, it is not necessary to generate all possible n-grams in each iteration. This method also ensures that the set of n-grams in the index is a prefix-minimal set, as the breadth-first search ensures that the shortest prefix of a useful n-gram is visited first. No n-grams in candidate sets of future iterations have the selected n-grams as prefixes, as the useful n-grams are never extended to generate a candidate n-gram.
FREE confines the search space by constraining the length $n$ of the n-grams selected. Briefly, there are two parameters to tune: 1) selectivity threshold $c$ that distinguish useful n-grams from useless n-grams, and 2) n-gram size $n$ to control the length of the index keys. In the original paper, the authors uses $c=0.1$ and $2\leq n\leq 10$ in their experiments. The accompanied index data structure used is inverted index with n-grams as keys and posting lists as values.
(a) Literal based query plan before checking index keys.
(b) Final query plan after index scan.
Figure 1. Example index search plan tree for regex <ahref=("|’).*ZZZ\. pdf("|’)> built by FREE, where the n-grams ", , Z, and pdf are indexed.
4.1.2. Regex Compilation
In order to also handle regex queries literals in alternative strings, FREE include a simple regex query plan compiler that compiles the overall index lookup to a plan tree with only logical AND and OR operators. Specifically, for a given regex, the compiler identifies literals on the primary level and literals, and then build a tree with the literals. FREE then checks if each of the literals or parts of the literals are indexed in the index, removing subtrees with no literals indexed, and substituting literal nodes with index keys, which are n-grams in the index.
We provide an example index search query plan in Figure 1 for the example regex, <ahref=("|’).*ZZZ\. pdf("|’)>, assuming that only ", , Z, and pdf are indexed. The literal tree shown in Figure 1(a) is built sorely by extracting all literals from the regex. Referring to the index, we notice that <ahref= and > are not indexed, and two n-grams in the literal ZZZ.pdf, Z and pdf are indexed. We simplify the plan to get Figure 1(b).
The evaluation of the index search plan tree then is a depth-first evaluation, where AND is a set intersection operation of two posting lists and OR is a set union operation of two posting lists.
4.1.3. Complexity Analysis
Let us denote the candidate set of each step $i$ as $M_{i}$. In the initial step, FREE scans the entire dataset to find all unique uni-grams and calculates their selectivities. The space needed is a temporary hashmap with the key being all unique unigrams of size $|M_{1}|=|\Sigma|$. The time complexity for the base step is $O(|D|+|\Sigma|)$. After scanning through the dataset, a pass is made on all unique unigrams to calculate their selectivity and move the useful unigrams from the candidate set into the index. In the next iteration, the remaining useless uni-grams in the candidate set of the previous step are extended by one character to construct a candidate set of bi-grams, $M_{2}$. The runtime space overhead of this iteration is $O(|M_{2}|)$, and the compute time overhead is $O(|D|+|M_{2}|)$.
For each iteration $i$, we have the runtime space overhead $O(|M_{i}|)$ and compute overhead $O(|D|+|M_{i}|)$. Since each step $i$ selects all useful n-grams of size $i$, and the temporary hashmaps of previous iterations are never revisited after candidate set $i$ is generated, for a FREE n-gram selection with $n\leq k$ for some integer $k$, we have the runtime space overhead $O(|M_{k}|)$ and compute overhead $O(k\cdot(|D|+|M_{k}|))$ for each iteration. As the number of n-grams of any size on a dataset is at most the number of characters in the dataset, we have the overall runtime space overhead $O(|D|)$ and the compute overhead as $O(2\cdot|D|)$ for all iterations.
4.2. BEST
Although FREE is effective in selecting n-grams with high filtering power for indexing, it has the limitation of not considering the query set. As a result, it cannot guarantee that the index will be as helpful when the character frequency distribution for the query literals differs from that of the dataset. BEST remedies this problem.
4.2.1. Selection Strategy
BEST utilizes both the dataset and the query set as sources for n-gram selection. In addition to considering the selectivity of n-grams within the dataset, it also takes into account the frequency of n-gram occurrences in the query set. This approach helps avoid selecting n-grams that do not benefit any regex query, as illustrated in the example in Section 3.
Assumption BEST-1 0.
It is beneficial to index an n-gram that does not appear in a data record, so that it can filter out the data earlier; it is beneficial to index an n-gram that appears in a query, so that it can be used to filter out data records for the query.
Assumption BEST-2 0.
The benefit of filtering out a data record $d_{i}$ is similar to the benefit of filtering out a data record $d_{j}$, where $i\neq j$.
With these assumptions, BEST abstracts the n-gram selection problem into a graph cover problem. Each data record, query, and n-gram is regarded as an individual node in the universe $U$. If an n-gram $g$ is present in a query $q$, there is an edge between $g$ and $q$; if $g$ is absent from an input string $d$, there is an edge between $g$ and $d$. In the context of our regex workload, each regex is matched once, so all edges have equal weights due to Assumption Assumption BEST-2. Each subset in $U$ represents a set of connected nodes interlinked by a single n-gram. Formally, we have the following construction.
Definition 4.3 (BEST).
Cover For a workload $W=(Q,D)$, the cover of an n-gram $g$ is the set of data records $d\in D$ and query $q\in Q$ pairs defined as follows:
$cover(g)=\{(q,d)\quad|\quad g\in q\land g\notin d,\quad\forall(q,d)\in Q\times D\}$
For a set of n-grams $G$, the cover of the n-gram set is the union of the cover of each n-gram in the set.
$cover(G)=\bigcup_{g\in G}cover(g)$
The n-gram selection problem is then transformed to a budgeted maximum set cover problem (Khuller et al., 1999) that aims to find the minimum set $G$ that achieves maximum cover. Formally, BEST define the value of indexing an n-gram with respective to the entire workload by its benefit.
Definition 4.4 (BEST).
Benefit For a workload $W=(Q,D)$ and an n-gram set $I$, the benefit of an n-gram $g\notin I$ is the number of additional query-data pairs covered by $g$ that is not already covered by $I$.
$benefit(g,I)=|cover(I\cup\{g\})-cover(I)|$
Definition 4.5 (BEST).
Cost For a workload $W$ consists of dataset $D$ and query set $Q$, the cost to index an n-gram $g$ is the storage overhead of $g$.
The cost is dependent of the index structure. In the original BEST paper, B+-tree is used, and the cost of $g$ is the number of leaf pointers corresponding to $g$, which number of data records in $D$ that contains $g$. For inverted index, the cost of $g$ is the size of its list, which is of size $s_{D}(g)$.
Definition 4.6 (BEST).
Utility For a workload $W$ consists of dataset $D$ and query set $Q$ and an existing n-gram set $I$ on $W$, the utility of indexing an additional n-gram $g$ is the ratio of its benefit over cost.
$utility(g)=\frac{benefit(g,I)}{cost(g)}$
BEST selects n-grams based on their utility as defined above. The brute-force method would be to iteratively select the n-gram with the highest utility, $g_{max}$, based on the workload $W$ and the set of already-selected n-grams $I$, among all possible n-grams with a positive benefit.
Assumption BEST-3 0.
The average selectivity of candidate n-grams is low for the data records and the literals in the queries.
Assumption BEST-4 0.
The average number of characters in all literals of a regex is much smaller than that of data records. The number of regex queries is much smaller than number of data records in the same workload.
Therefore, according to the sparsity assumption Assumption BEST-3, BEST choose to use adjacency lists: Q-G-list and G-D-list rather than two matrices of sizes $|Q|\cdot|G|$ and $|G|\cdot|D|$ to store the existence of n-grams in the workload to reduce space usage. The candidate n-grams, queries, and data records are each assigned an index number. The indices of Q-G-list correspond to query numbers, and each element is a list of n-gram numbers in this query. Similarly, the indices of G-D-list correspond to n-gram numbers, and each element is a list of data record numbers which the n-gram is in.
4.2.2. Approximation Techniques
The budgeted maximum set cover problem optimization problem is NP-hard (Lund and Yannakakis, 1994; Khuller et al., 1999). Besides, the search space is the product of number of possible query-data pairs, $|Q|\cdot|D|$ and the number of possible n-grams, $|G|$. By Assumption Assumption BEST-4, we assume $G=G(Q)$. When the query size and the dataset size is large, the number of candidate n-grams can be prohibitively large to select the n-grams using brute-force method. BEST employs several techniques to get an approximate result.
Pruning. BEST introduce pruning of n-grams by selectivity to reduce the number of the candidate n-grams. With the same assumption as Assumption Assumption FREE-1, BEST prune n-grams that has selectivity larger than a threshold $c$.
Parallelism by Clustering. BEST clusters the regex queries into small groups of similar queries that contains largely overlapping n-grams. Within each small group $Q_{i}$, the computation will be $O(|Q_{i}|\cdot|D|\cdot|G(Q_{i})|$. Formally, the distance is calculated as
$Dist(q_{1},q_{2})=\frac{|(G(q_{1})-G(q_{2}))\cup(G(q_{2})-G(q_{1}))|}{|G(q_{1})\cap G(q_{2})|}$
Since queries with similar set of n-grams are clustered together, we minimize the number of n-grams for each subproblem, that is, the size of $G(Q_{i})$. The clustering allows BEST to divide the search in each iteration into smaller sub-problems that allows for parallel computation. The intermediate cost and benefit results from all sub-problems are aggregated at the end of each iteration to select the n-gram with maximum utility.
Workload Reduction. When the size of the query set $Q$ is large, BEST selects a representative sample $Q^{\prime}\subseteq Q$ to further reduce the computation overhead. To ensure that the sample is representative, BEST use the same clustering technique to cluster the literals in $Q$ into clusters. When the clusters stabilize, the median query of each group is selected into $Q^{\prime}$.
4.2.3. Complexity Analysis
First, the workload reduction technique is applied, reducing the query set $Q$ to a reduced set $|Q^{\prime}|=\frac{|Q|}{t}$ for some $t\geq 1$ and the candidate n-gram set being $G^{\prime}$. It is reasonable to assume that $|G^{\prime}|\simeq\frac{|G|}{t}$. Let $\mu$ be the average number of n-grams in a query and $T$ be the maximum number of iterations of the k-median algorithm. Then, the workload reduction time is $O(|Q|\cdot|Q^{\prime}|\cdot\mu\cdot T)$. We use a suffix tree to efficiently enumerate all possible n-grams in the queries, which will take $O(|G^{\prime}|)$ time to build, enumerate, and insert them to a hash-map. A similar clustering is then carried out for creating sub-problems for parallelism. Empirically, according to the original BEST paper, dividing the problem to sub-problems reduces both time and space complexity by approximately $5$ even if we run it on a single thread. The use of the two adjacency lists Q-G-list and G-D-list rather than two matrices reduces the space overhead for utility calculation from $O(|G|\cdot(|D|+|Q|))$ to $O(\mu\cdot(|D|+|Q|))$. By Assumption Assumption BEST-4, $\mu\ll|G|$, and therefore, the memory overhead during computation is small. Note that we still need $O(|G|\cdot(|D|+|Q|))$ time to build both data structures. The two adjacency lists contributes to the majority of extra space overhead for BEST. In each iteration, BEST will examine all remaining candidate n-grams and their benefit considering the index $I$ built so far. For each pair $(q,d)\in(Q^{\prime}\times D)$, it checks if the pair covered by each candidate n-gram $g$. The time complexity of each iteration is $O(|G^{\prime}|\cdot|Q^{\prime}|\cdot|D|)$. The overall time complexity for BEST including workload reduction and clustering can be expressed as $O(|Q|\cdot|Q^{\prime}|\cdot\mu\cdot T+|G|\cdot(|D|+|Q|)+\frac{|I|\cdot|G^{\prime}|\cdot|Q^{\prime}|\cdot|D|}{5})$, which simplifies to $O(\frac{|I|}{5t^{2}}\cdot|G|\cdot|Q|\cdot|D|)$.
4.3. LPMS
Despite several optimizations in the BEST algorithm, its complexity analysis is still too large. LPMS remedies this issue by introducing approximations in the algorithm for BEST via integer programming.
4.3.1. Selection Strategy
Similar to BEST, LPMS also uses both the query set and the dataset as sources of n-gram selection. LPMS also incorporates Assumption Assumption FREE-1 that an n-gram that eliminates more data records is more useful, but it incorporates the impact of queries by adjusting it with the selectivity of n-grams in queries and the length of the n-gram. Formally,
Definition 4.7 (LPMS).
Coverage The coverage of an n-gram $g$ is defined as the ratio of the support of $g$ in the dataset $D$ and the support of $g$ in the query set $Q$, normalized by the n-gram length.
$cv(g)=\frac{s_{D}(g)}{|g|\cdot s_{Q}(g)}$
Similar to Assumption Assumption BEST-4, we assume $G=G(Q)$. Using the binary variable $x_{g}=\mathbbm{1}\left[g\in I\right]$, we form the objective function as $\sum_{g\in G}cv(g)x_{g}$. Note that $x_{g}\in\{0,1\}$ $\forall g\in G$. To provide good approximation to the optimal solution, the approximate set of n-grams should satisfy the following:
Assumption LPMS-1 0.
The index should filter out at least as many data records compared to any candidate n-gram.
Let $g_{j}$ to represent the $j$-th n-gram in the candidate n-gram set $G$ and $q_{i}$ to represent the $i$-th query in the query set $Q$. LPMS constructs a matrix $A$ of size $|Q|\times|G|$ and a vector $b$ of size $|Q|$ for constraint calculation, where $A_{i,j}=s_{D}(g_{j})\cdot\mathbf{1}{g_{j}\in G(q_{i})}$ and $b_{i}=\min_{g\in G(q_{i})}{s_{D}(g)}$. This setup allows us to establish the constraint of the integer program, formalizing the Assumption Assumption LPMS-1 into the constraint $Ax\geq b$.
However, the search space for all possible n-grams $G$ remains too large when the query set is large. LPMS adopts an iterative approach to select a prefix-minimal n-gram set from FREE. In the $i$-th iteration, LPMS generates the candidate n-gram set $G_{i}$ with n-grams of size $i$ from all the useless n-grams from $G_{i-1}$. After solving the integer program in the $i$-th iteration, we insert the set of n-grams $I_{i}$ with $x_{g}=1$ for all $g\in I_{i}$ into the index, and the remaining n-grams $G_{i}\setminus I_{i}$ are used to extend and generate $G_{i+1}$.
Solving the integer program is challenging, as the search space is $|G_{i}|^{O(|G_{i}|)}$ (Lenstra, 1983). LPMS approximates the problem using linear programming with relaxation, as follows:
minimize $\sum_{g\in G}cv(g)x_{g}$
subject to $Ax\geq b$
$0\leq x_{g}\leq 1\quad\forall g\in G$
4.3.2. Complexity Analysis
To calculate the space overhead, let’s examine the sizes of each component. Both coverage and the output of the linear program are vectors of size $|G_{i}|$. As previously discussed, $A$ has a size of $|Q|\cdot|G_{i}|$ and $b$ has a size of $|Q|$. Summing these, the space overhead for LPMS n-gram selection is $O(|Q|\cdot|G_{i}|)$. During the algorithm runtime, coverage is calculated using $s_{D}(g)$ and $s_{Q}(g)$. By utilizing coverage for each n-gram rather than cover, LPMS reduces the time complexity for constructing the coverage vector to $O(|G_{i}|\cdot(|D|+|Q|))$ for each iteration. The linear program runs in polynomial time $O(|G_{i}|^{2.5})$ (Vaidya, 1989). In practice, the size of the index key typically does not exceed 10, a number also used as the upper bound for n-gram size in FREE. Therefore, we can consider the small number of iterations as a constant, making the overall computational complexity of LPMS $O(\sum_{i\in[10]}|G_{i}|^{2.5}+|G_{i}|\cdot(|D|+|Q|))$.
5. Experiment Setup
Due to the absence of source code from the papers, we implemented the three n-gram selection methods ourselves. For each method, we also developed their corresponding query parsers (if any) and index structures.
5.1. Benchmark Framework
The benchmark framework is designed to facilitate the comparison of n-gram selection techniques (FREE, BEST, and LPMS) for regex indexing, with a focus on modularity, extensibility, and reproducibility across a range of workloads. We summarize its detailed architecture and workflow in Figure 2.
The end-to-end process follows a seven-step sequence, as illustrated in the framework diagram Figure 2. First, the user specifies inputs (dataset, regex queries, method name, and configurations), which are then processed by the framework. The selected n-gram strategy processes the workload, identifying optimal n-grams for index construction. These n-grams are used to build an index (e.g., inverted index or B+-tree). The index search plan compiles by examining the regex patterns and extracting the n-grams in their literals to create plans for index lookup, filtering candidate data points. The regex engine then verifies these candidate data points, discarding false positives, and generating metrics to quantify performance. Finally, the results are aggregated and exported. The framework operates in a pipeline architecture divided into three phases: input processing, index construction, and regex evaluation. The components of each phase are grouped within dashed squares in Figure 2.
Input Processing. The framework begins by accepting a text file containing the string dataset and a text file with the regex query set. Users also specify n-gram selection methods and their parameters, including n-gram length, selectivity thresholds, maximum number of n-grams, thread counts, and more for index building. Optionally, a new query set can be provided at runtime, enabling dynamic evaluation of the index against unseen query workload. These inputs are standardized to a unified format across experiments. During this phase, all data in the workload is loaded into memory.
Index Construction and Searching. In the index construction phase, one of the three n-gram selection strategies is applied to the workload. The selected n-grams are used to build an index, such as an inverted index or a B+-tree. This phase leverages multi-threaded execution if specified by the user. After the index is built, the index search plan is compiled if necessary. Each strategy has its own corresponding type of index structure and index search plan, as shown in the blue dashed bounding box in Figure 2. FREE and LPMS use inverted indexes, while BEST uses a B+-tree for the index structure. During index lookup, all n-grams in both the set of regex query literals and index keys are extracted. Only FREE requires an additional intermediate step to generate an index search tree.
Regex Evaluation and Result Output. The final phase processes all possible data points after index lookup and validates them using a regex engine (e.g., RE2, PCRE2) to perform exact matches and eliminate false positives. Metrics such as index construction time, runtime memory consumption, workload processing time, and false-positive rates are measured during index construction or after regex evaluation.
Figure 2. Benchmarking framework overview.
Table 2. Workload $W=(Q,D,\Sigma)$ statistics. We use $\bar{|d|}$ to denote average data record size (in bytes) and $\bar{TP}$ to denote the average number of actual matches of all queries. Synthetic has two query sets – the first set of 500 queries that is used to build the index, and a second set of 100 queries that is used to test the impact of queries that are not indexed.
Workload $|Q|$ $|D|$ ||$\sum$|| ∑ | $\bar{|d|}$ $\bar{TP}$ Dataset Size
Webpages 10 695,565 255 68,650 86,524 47 GB
DBLP 1000 305,798 122 38 914 32 MB
Prosite 101 111,788 22 416 722 45 MB
US-Acc 4 2,845,343 99 405 92,042 1.1 GB
SQL-Srvr 132 101,876,733 114 139 50,356 14 GB
Synthetic 500/100 5,000 16 32 628 163 KB
5.2. Workloads
We use several real-world text datasets and queries with varying numbers of queries, data strings, alphabet sizes, average string lengths, and average matches per query for a comprehensive analysis. We summarize the workload characteristics in Section 5.1. Among the workloads, three (Webpages, DBLP, and Prosite) are those used by the original FREE, BEST, and LPMS papers to evaluate their methods. The other two workloads (US-Accand SQL-Srvr) are more recent and feature larger datasets.
Webpages. This workload was used by FREE. In the original paper, the authors utilized 700,000 random web page HTML files downloaded in 1999, along with 10 regex queries suggested by researchers at IBM Almaden (Cho and Rajagopalan, 2002). While the exact query set is included in the paper, the dataset is not. Since we could not locate the original dataset, we constructed a similar dataset using web pages from 2013 stored in Common Crawl (Common Crawl Foundation, 2013). We chose 2013 data because it is relatively close to 1999, ensuring that most of the regexes constructed for the 1999 dataset would still have matches in the 2013 dataset. We selected the web pages to ensure a relatively balanced number of matches for each regex query.
DBLP. This workload was used by BEST. The authors collected 305,798 (Author-Name, Title-of-Publication) tuples as the dataset. We used the DBLP-Citation-network dataset (Tang et al., 2008) and selected the same number of entries uniformly at random. The query set is constructed by choosing author last names from the pool of author names uniformly at random to obtain $1000$ queries. We then constructed the regex query for each last name by appending .+ and a space in front of the last name.
Prosite. This workload was used by LPMS. Following the paper’s description, we selected 100,000 protein sequences from the PFAM-A database (Mistry et al., 2020) and chose 100 Prosite signatures (Sigrist et al., 2012), transforming them into 100 regular expressions.
US-Acc. This is an open-source dataset containing descriptions of traffic accidents in the United States from February 2016 to March 2019 (Moosavi et al., 2019). The dataset has 2,845,343 strings. The authors included four regex queries on the dataset in the original table.
SQL-Srvr. This is a production workload consisting of in total 101,876,733 log messages generated by Microsoft SQL Server and 132 regex queries used for data analysis tasks on this dataset.
Synthetic. To test the robustness of the n-gram selection methods in handling unseen queries, we build off the synthetic dataset in the original LPMS (Tsang and Chawla, 2011) paper. The dataset consists of 5000 strings each constructed with alphabet ‘A-P’. The string size follows a geometric distribution with $p=1/32$. The index building query set is generated by randomly selecting 10% from the dataset, and a random slice is used to create a regex lit1regexlit2. lit1 has 1-5 characters, lit2 has 0-5 characters, and regex matches any $m$ characters where $1\leq m\leq 50$. During the query phase, we generate regexes of a similar format from 2% of the data records.
5.3. Metrics
We compare the three methods on the following aspects:
N-Gram Index Construction Time ($T_{I}$). We measure the time required to select all n-grams for indexing and building the index, denoted by $T_{I}$, after loading the necessary data into memory. This aspect is crucial, as we deal with much larger datasets than when these n-gram selection algorithms were originally presented. The selection time may become prohibitively long, rendering the methods impractical.
Precision. We use micro-average precision to compare the filtering power of an index. This metric best describes the effectiveness of the selected n-grams and is commonly used in information retrieval (Zhai and Lafferty, 2001; Liu et al., 2024). It provides a balanced measure across all queries with different numbers of matches, without giving disproportionate weight to any individual query. It aggregates actual data records that match the regex query (true positives, denoted by $TP$) and data records that pass the index filtering but do not match the regex query (false positives, denoted by $FP$). For a workload $W=(Q,D)$, the overall precision on index $I$ is:
$Prec_{W}=\frac{\sum_{q\in Q}\#TP_{q}}{\sum_{q\in Q}\#TP_{q}+\sum_{q\in Q}\#FP_{q}}$
Workload Matching Time ($T_{Q}$). We measure the time to run the workload, denoted by $T_{Q}$, to demonstrate the runtime gains provided by the index.
Runtime Space Usage ($S_{Q}$). This metric is the peak space used when running a experiment.
Index Size ($S_{I}$). We measure the size of the index constructed, which is the total size of the n-gram keys, the posting lists, and necessary index data structure (such as the B+-tree).
5.4. Hardware and Implementation Details
We implemented FREE and its query parser, integrating it with an inverted index. For LPMS, we also used the same inverted index structure to ensure consistency. We implemented BEST with a B+-tree as the supporting index data structure. We used Gurobi Optimizer (Gurobi Optimization, LLC, 2024) as the linear program solver. All experiments were conducted on an Azure Standard_E32-16ds_v5 machine with an Intel(R) Xeon(R) Platinum 8370C CPU @ 2.80GHz, 16 vCPUs, 256 GB of memory, and 1 TB of disk storage. Our benchmarking framework is written in C++17 and compiled with the -O3 flag. Google’s RE2 (release version 2022-06-01) (Google, [n.d.]) as the regex engine for regex evaluation.
Each of the three methods has different configurable parameters. For BEST and FREE, we vary the selectivity threshold $c$ by picking values between 0.01 and 0.7. For FREE, we also vary the upper bound of n-gram length, $\max_{n}$, among values: 2, 4, 6, 8, and 10. Another parameter we vary in FREE is whether the selected n-gram set is prefix minimal or pre-suf minimal. We set a hard threshold of 3 hours for all experimental runs. If BEST fails to finish within this time limit, we reduce the workload by a factor of 0.1%, 0.2%, 0.5%, and 0.85%. LPMS allows us to configure whether to use deterministic relaxation (LPMS-D) or random relaxation (LPMS-R) of the linear programs. Similarly, we set a maximum number of n-grams for LPMS in case a workload takes too long to run. For all configurations of all methods, we use 16 threads. To fairly assess the methods, we compare the resulting indexes with a similar number of n-grams selected. Since FREE and LPMS do not natively support dynamic constraints on the index size, we implement an early-stopping mechanism for all three methods. By providing an optional parameter of max number of keys, the n-gram selection methods will stop once the limit is reached. We choose the values of max number of keys on a case-by-case basis for each workload, as specified in Section 6.1.
6. Experiments
In this section, we run the aforementioned workloads on the three indexes with varying parameters. In particular, we aim to answer the following questions:
  1. Q.1
    How does each method perform on different datasets?
  2. Q.2
    How do the characteristics of each workload affect the usefulness of n-gram selection methods?
6.1. Query Performance
We determine the effectiveness of the index by evaluating the filtering effectiveness of the selected set of n-grams. The higher the precision of an index, the more effective the set of n-grams selected. Each configuration of a selection method generates a specific set of n-grams of varying sizes. To ensure a fair comparison, we compare selected sets of similar sizes, as indexing all possible n-grams would yield the highest precision for the workload. For each n-gram selection method under a specific key number constraint $K$ $s.t.$ $|I|\leq K$, we select the configuration with the highest precision. Note that throughout the section, we report the numbers for space usage and query performance for the configuration that achieves the best possible precision given a constraint on $K$.
6.1.1. Workload DBLP
Taking the DBLP workload as an example, we run BEST, FREE, and LPMS with varying parameters. After obtaining the results, we compare the performance of each n-gram selection method based on the upper bounds of the number of n-grams (i.e. index entries in $I$). Specifically, for this workload, we examine indices with a number of keys is at most $5000$. We summarize the results in Section 6.1.1, where the metric columns are variables introduced in Section 5.3.
Table 3. Index cost and query performance on DBLP.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
150 BEST 19.7 533 2.194 33.433 0.235
FREE 25.3 2 0.171 37.299 0.051
LPMS 19.8 3 0.155 34.028 0.048
500 BEST 15.2 8762 2.670 54.306 0.242
FREE 20.6 2 0.242 66.774 0.165
LPMS 17.7 6 0.158 40.657 0.070
1000 BEST 15.2 8762 2.670 54.306 0.242
FREE 18.7 1 0.287 75.354 0.219
LPMS 17.0 808 1.309 41.499 0.071
2000 BEST 15.2 8762 2.670 54.306 0.242
FREE 17.4 1 0.264 68.442 0.321
LPMS 17.0 808 1.309 41.499 0.071
3000 BEST 15.2 8762 2.670 54.306 0.242
FREE 16.9 2 0.304 83.069 0.524
LPMS 17.0 808 1.309 41.499 0.071
4000 BEST 15.2 8762 2.670 54.306 0.242
FREE 16.9 2 0.304 83.069 0.524
LPMS 17.0 808 1.309 41.499 0.071
5000 BEST 15.2 8762 2.670 54.306 0.242
FREE 16.6 2 0.316 85.266 0.535
LPMS 17.0 808 1.309 41.499 0.071
[…]
When relaxing the index size constraint to 1000 n-grams, FREE ($\max_{n}=4$,$c=0.5$) generated an index of size 810 that achieved a precision of 0.219, slightly lower than BEST($c=0.5$). However, this configuration has a construction time of around 1 second, which is several orders of magnitude lower than both BEST and LPMS-D. FREE also consumes significantly less memory during computation. It uses 0.287 GB, whereas LPMS requires 1.3 GB and BEST 2.67 GB of memory to store the data and perform computation. For LPMS, with more n-grams selected, its second configuration achieves 10 times less index construction time and higher precision compared to using a smaller number of keys, resulting in a workload evaluation performance of 17.1 seconds, which is similar to the other methods. Although LPMS selects 929 keys to index, which is larger than the 810-key inverted index by FREE, it still has a smaller index size. This again demonstrates that LPMS tends to select n-grams with low frequencies in the dataset.
Figure 3. Index sizes of indices with different key numbers built by the three methods on DBLP workload.
As the number of n-grams allowed in the index increases, starting from 2000, more indices from different configurations of FREE fall into this range. These configurations have low selectivity thresholds ($c<0.15$) and $n=2$ compared to the best index generated by FREE when $K=1000$, selecting n-grams that do not cover enough strings. This is demonstrated by the smaller sizes of these indices in Figure 3 even as number of keys increase, representing their short posting lists. The best configurations of FREE($\max_{n}=2$,$c=0.2$) achieve higher precision than BEST. The precision increases from 0.321 at 2000 keys to 0.535 at 5000 keys upper limit by FREE($\max_{n}=4$,$c=0.1$), without a significant increase in construction overhead. As expected, the index size grows with the number of keys. However, the improvement in workload performance is less significant compared to the change in precision. In the DBLP workload, each string in the dataset is short, resulting in very quick regex evaluation.
Despite its precision, BEST achieves the best workload query performance. The index built by BEST has the smallest number of index keys, totaling 363. In comparison, a similar precision index by FREE requires 1440 keys. A larger index incurs greater index lookup overhead, which is significant when the dataset consists of short strings that match quickly after passing the filter. Another source of overhead for FREE comes from traversing its query plan trees, which are optimized for regexes with alternation (logical OR).
{customframe}
Insights
FREE performs best on the DBLP workload. Generally, if the workload has a large number of queries that are not skewed (i.e., all very similar and covering a small subset of the dataset) and the strings in the dataset are not long, it is best to choose FREE. When the query set is large, BEST and LPMS require significant computation time and memory, owing to their higher time and space complexities. Additionally, for a large-sized and balanced query set, frequent n-grams in the datasets are likely to be covered in the queries as well. Therefore, FREE benefits from these two aspects and outperforms the other two methods.
6.1.2. Workload Webpages
Next, we examine a workload that is drastically different from the DBLP workload. As shown in Section 5.1, Webpages has a significantly smaller number of queries in its workload compared to DBLP. Each string in the Webpages dataset is an HTML file for a webpage, making the average string length the longest among all the workloads. We run the workloads using the three methods and summarize the experimental results in Section 6.1.2.
We observed that the computational overhead of BEST is significantly reduced compared to DBLP due to the smaller query set. BEST achieves reasonable precision for performance improvement with a very small index size. However, since the query set is not representative, this small index may not benefit on unseen queries.
Looking at the first row of Section 6.1.2, by selecting only 3 to 4 n-grams for indexing with BEST($c=0.1$), the index achieves a precision of 0.138. By manually early stopping FREE($\max_{n}=2$,$c=0.02$) at 5 n-grams selected, it achieves comparable precision of 0.123, while incurring around 20$\times$ lower time overhead.
Table 4. Index cost and query performance on Webpages.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ GB $S_{I}$ MB $Prec$
5 BEST 422 8587 150 0.1 0.138
FREE 532 428 148 0.1 0.123
$\mathbf{1.7\times 10^{5}}$ BEST 422 8587 150 0.1 0.138
FREE 353 1184 159 357.9 0.163
$\mathbf{3.6\times 10^{5}}$ BEST 422 8587 150 0.1 0.138
FREE 240 5965 240 2007.4 0.301
$\mathbf{1.5\times 10^{6}}$ BEST 422 8587 150 0.1 0.138
FREE 239 7480 233 524.9 0.302
To attain a similar precision without early stopping, FREE($\max_{n}=2$,$c=0.5$) requires an index with more than 10000 times the number of keys and more than 3000 times the index size at $K=1.7\times 10^{5}$ The difference in the number of keys results from the different sources of n-grams considered during selection for the two methods. BEST chooses index keys from the intersection of n-grams in both the query set and the dataset, and a smaller number of queries means a smaller number of potential n-grams, effectively constraining the index size and computation overhead. On the other hand, FREE derives its index keys from the dataset, which contains many more distinct n-grams, leading to the selection of a large set of n-grams that meet the requirements defined by the specific configuration. Looking at the precision, we see that FREE, with a 0.025 higher precision than BEST, achieves a 1.2×1.2$\times$1.2 × overall query time speedup, reducing the time from 421.6 to 353.0 seconds.
When the number of keys allowed $K$ increases from $1.7\times 10^{5}$ to $3.6\times 10^{5}$ , the precision of the index by FREE($\max_{n}=2$,$c=0.7$) increases significantly from 0.163 to 0.301. The index construction time also increases significantly. The query time of the best index built by FREE for this $K=3.6\times 10^{6}$ is 1.47×1.47$\times$1.47 × faster and 1.75×1.75$\times$1.75 × faster than the index built with the three n-grams selected by BEST. Compared to the DBLP workload, the precision increase in Webpages corresponds to a more significant workload runtime decrease from 353 seconds to 240 seconds. In this workload, each webpage is long, and the cost of regex evaluation on one string is high. Therefore, eliminating more strings with the index significantly improves the performance. These advantages might not justify the large index size compared to BEST for this specific workload. However, it does make the index more robust to unseen queries on the same dataset. LPMS did not finish for this dataset within the set time frame due to the large average number of characters per line of the dataset. Unlike BEST where the data structures for intermediate calculation are adjacency lists due to the sparseness Assumption Assumption BEST-4, LPMS builds matrices of size $|Q|\times|G|$ as input to integer program solver, where the actual computational and space overhead is too large.
{customframe}
Insights
BEST is suitable for a workload like Webpages where the query set is small. BEST selects the set of n-grams that achieves near-optimal precision, with the optimality resulting from the long computation time. With a small query set, the index construction time is reasonable. However, for a workload where each document entry is large, the precision and robustness of the index become more important, especially if the dataset is likely to be queried with different regex queries.
6.1.3. Workload Prosite
Prosite has the largest query size to number of records in the dataset ratio among the real-world workloads. The strings in the dataset have a mean length of 416 characters. Two distinguishing characteristics of Prosite are: 1) it has the smallest alphabet size of 22, and 2) it has very short literal components in its the regex queries.
Table 5. Index cost and query performance on Prosite.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
50 BEST 141.5 400 0.931 9.4 0.00826
FREE 154.0 2 0.227 1.5 0.00651
LPMS 151.8 14 0.263 26.2 0.00708
100 BEST 139.7 856 0.967 14.9 0.0089
FREE 153.6 2 0.242 5.7 0.00652
LPMS 151.8 14 0.263 26.2 0.00708
500 BEST 139.7 856 0.967 14.9 0.00890
FREE 150.8 3 0.595 141.4 0.00687
LPMS 151.8 14 0.263 26.2 0.00708
Section 6.1.3 shows the results. BEST performs the best for this workload, generating index with highest precision while using a small number of keys of 364. The small alphabet size and small literal size per query make the number of possible n-grams considered in each calculation iteration small, and thus drastically reduce the index construction runtime overhead, which is the biggest advantage of BEST in other workloads.
When indexing with 50 keys, FREE($\max_{n}=2$,$c=0.15$) generates an index with a higher precision of 0.00651, while taking the least computational overhead and smallest index size. LPMS-D, taking slightly longer to select the n-grams and generate the index, achieves a precision of 0.00708 within 14 seconds. It achieves similar query performance compared to the index by FREE, although the index size is significantly larger, as it selects more frequent n-grams as keys. BEST($c=0.7$) achieves the highest precision, 0.00826, among the three methods. Its query time is also more than 10 seconds lower than the other two methods. However, it requires 29×29$\times$29 × the index construction time compared to LPMS and 200×200$\times$200 × compared to FREE. It also consumes significantly more compute memory, nearly 3×3$\times$3 × higher than that of FREE and LPMS when $K=50$ and $K=100$. Overall index precisions are low for Prosite workload, as the average length of literals is very short in the regexes.
When $K=100$, the index by FREE($\max_{n}=2$,$c=0.2$) has almost identical precision with the 50-key index constructed by the same method. The workload running time for LPMS-D is slightly faster than FREE, with a similar compute memory. Index by BEST($c=0.7$) has a higher precision of 0.0089, resulting in a more than 10 seconds workload performance improvement over other methods. BEST takes 856 seconds to construct the index, much longer than other methods, but having smaller index size than LPMS. As the key size constraint loosens to $K=500$, FREE selects an index which only marginally decreases its overall workload running time from 153.6 seconds to 150.8 seconds, while using 395 keys (330 more keys than when $K=100$). The size of the index increases to 24.6×24.6$\times$24.6 × compared to the index built by FREE when $K=100$, and is much larger than the index sizes of other methods. However, its precision is still less than that of BEST and LPMS with a much smaller number of keys.
The results align with the characteristics of the methods: BEST trades off n-gram selection time for higher precision n-grams tailored to the workload, while LPMS approximates the results of BEST by trading off some precision for lower index construction overhead. When comparing the index construction memory and time overhead to the results for the same $K$ in DBLP, we observe that for Prosite, BEST has significantly smaller overhead that is not proportional to the dataset size difference. In Prosite, the possible literal set is small due to the limited alphabet size and short query literal sizes. Consequently, the computation time and space are significantly reduced in each step when BEST needs to count the number of lines for each n-gram.
{customframe}
Insights
For workloads where query literals are short and/or the alphabet size is small, BEST can select an n-gram set with high filtering precision while incurring reasonable computational and storage overhead. This advantage arises from the decreased number of potential n-grams to consider. FREE does not benefit from this advantage since it only looks at n-grams in the dataset. LPMS benefits less from the characteristics of the workload due to the fixed overhead for constructing the integer program.
6.1.4. Workload US-Acc
This workload contains the smallest number of queries. Compared to the Webpages workload, which also has a small query set, the datasets and queries are generated by a limited number of templates. The data records in the dataset consist of the location description (e.g., ”At I-270, Between OH-48/Exit 29 and Dayton Intl Airport Rd/Exit 32”) and a brief description of the accident. Each string in the US-Acc dataset is much shorter.
Table 6. Index cost and query performance on US-Acc.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
30 BEST 1.451 63 2.204 84.80 0.510
FREE 4.940 29 0.969 0.02 0.032
LPMS 1.353 235 2.876 60.85 0.572
1000 BEST 1.451 63 2.204 84.80 0.510
FREE 1.708 31 3.235 1003.99 0.460
LPMS 1.353 235 2.876 60.85 0.572
5000 BEST 1.451 63 2.204 84.80 0.510
FREE 1.297 46 3.199 1070.97 0.811
LPMS 1.353 235 2.876 60.85 0.572
Section 6.1.4 shows the results. LPMS-D performs the best for the US-Acc workload, achieving a precision of 0.572 using only 12 keys. Although it takes 3.7×3.7$\times$3.7 × the index computation time compared to BEST($c=0.7$) and slightly more computation space overhead, LPMS uses 0.4×0.4$\times$0.4 × less index space to achieve this higher precision. When $K$ increases to 1000, FREE generates an index with 837 keys and a precision of 0.46, which is much lower than BEST and LPMS with fewer than 30 keys. This is because FREE selects n-grams based solely on their selectivity in the dataset, omitting those n-grams that are frequent in both the regex query and the dataset. In fact, the configuration of FREE that generates the highest precision with fewer than 1000 keys uses a selectivity threshold $c=0.7$, the largest across all of our experiments.
{customframe}
Insights
For workloads where the query literals are very short and common in the dataset, LPMS can quickly locate a small set of n-grams with high precision. Both BEST and FREE remove n-grams that are frequent in the dataset, as they both adhere to Assumption Assumption FREE-1 that does not hold for the dataset.
Table 7. Index cost and query performance on SQL-Srvr.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
50 FREE 3457 1342 76.7 6643 0.000945
LPMS 1421 4569 147.5 4474 0.009855
5000 FREE 1473 1483 74.0 6643 0.002698
LPMS 1421 4569 147.5 4474 0.009855
6.1.5. Workload SQL-Srvr
The SQL-Srvr workload has dataset characteristics similar to the US-Acc workload, as it is generated from several formatted strings—in this case, system log reporting strings. We summarize the results in Section 6.1.4. The size of both the query set and dataset are much larger than those in US-Acc, and therefore, BEST failed to finish within the set time-frame.
LPMS performs better than FREE for similar reasons: it considers the query set without overlooking common n-grams. When we limit the number of n-grams to 50, LPMS-D selects an n-gram set with a precision of $9.855\times 10^{-3}$, which is 10×10$\times$10 × higher than the index generated by FREE($\max_{n}=2$,$c=0.7$) with the same number of keys. This is because FREE blindly selects selective n-grams without considering the queries. Given the variability in the values of the formatted system logs (e.g., VM IDs, cluster names), FREE may index short n-grams from them, resulting in an index that hardly benefits the actual query set. Consequently, the query time for LPMS is 2.43×2.43$\times$2.43 × faster. Although LPMS takes 3.4×3.4$\times$3.4 × longer and requires 1.9×1.9$\times$1.9 × more construction space and time overhead, the index built by LPMS uses 1.48×1.48$\times$1.48 × less space. Considering that LPMS and FREE use the same inverted index implementation and they have the same number of keys (50), the only difference in their index sizes comes from the posting lists lengths, which are the IDs of the data records. This suggests that LPMS selects n-grams that are more selective in the dataset. Combined with its higher precision and faster query time, LPMS selects n-grams that are both more selective in the dataset and more common in the query set, making them more beneficial. The results remain consistent until the upper limit of the number of n-grams grows to 5000. At this point, FREE($\max_{n}=2$,$c=0.03$) selects an n-gram set of size 4643, increasing the precision to $2.68\times 10^{-3}$. This is achieved with only a slight increase in index construction overhead and nearly identical index size compared to indexing with 50 n-grams selected by LPMS. However, the precision and query time are still worse than those of LPMS with 50 n-grams.
{customframe}
Insights
When query literals are common, the n-grams that have high benefit may also have high selectivity, which contradicts Assumption Assumption FREE-1. For workloads where query literals are common in the dataset, even when the query literals are long, LPMS selects the n-gram set that improves query matching time the most. This is due to LPMS’s design choice of not discarding any n-grams based on selectivity threshold.
Table 8. Index cost and query performance on Synthetic.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
20 BEST 0.829 3.398 0.042 0.091 0.2084
FREE 0.190 0.027 0.016 0.446 0.2672
LPMS 1.211 3.720 0.053 0.010 0.1479
100 BEST 0.444 31.659 0.043 0.452 0.4793
FREE 0.190 0.027 0.016 0.446 0.2672
LPMS 1.063 3.791 0.052 0.016 0.1757
300 BEST 0.061 90.661 0.043 1.125 0.6451
FREE 0.064 0.032 0.015 1.122 0.6453
LPMS 1.063 3.791 0.052 0.016 0.1757
6.1.6. Robustness Test
For this test, we use the synthetic workload with different sets of regexes for index construction and query matching to test the robustness of each method for unseen queries. We summarize the result in Section 6.1.5.
For $K=20$, FREE($\max_{n}=2$, $c=0.7$) builds an index with 16 keys that achieves the highest precision of 0.267. It also has the lowest index construction overhead. BEST($c=0.2$) achieves slightly lower precision of 0.208, but the query time is $\times4$ more than that of FREE. LPMS-D performs the worst with the lowest precision and highest query time. Note that BEST and LPMS also have significantly higher computational overhead compared to FREE.
As $K$ increases to 100, BEST($c=0.5$) achieves the highest precision. It’s index construction time also increases by $\approx$3 times to 31.7s. This is more than 1000×1000$\times$1000 × longer than FREE. Note that the query times for BEST are not the lowest among the three methods, although it has the highest precision. This is due to the non-uniformness in the data record lengths, which breaks Assumption Assumption BEST-2.
As we increase to $K=300$, FREE($max_{n}=2$, $c=0.12$) achieves the highest precision of 0.6453. BEST($c=0.2$) achieves similar precision and similar overall query time compared to FREE, but with a much larger index construction time which is 2833×2833$\times$2833 × slower than FREE. In all cases, LPMS performs unfavorable to the other two methods, since the n-gram selection strategy of LPMS is solely based on the benefit calculation using the query set and the dataset, without any heuristic like the pruning step of BEST.
{customframe}
Insights
For scenarios when a unseen queries on a fixed dataset are expected, the strategy that takes advantage of the query set (e.g. BEST and LPMS) may fall short. FREE, selecting n-grams based on the dataset, is robust in this setting.
7. Learnings and Future Work
Figure 4. N-gram selection and regex index method decision tree by workload characteristics.
The evaluation shows that the optimal choice of n-gram selection strategies among FREE, BEST, and LPMS is strongly influenced by workload characteristics (such as the query set size, query literal sizes, alphabet size, data patterns, etc.). Since no single strategy suits all workloads, this analysis emphasizes the importance of selecting methods based on workload characteristics and provides a general guideline. We summarize the insights in the decision flow chart in Figure 4 as a guideline for practitioners. This guideline recommends 1) choosing FREE for large and diverse workloads or when unseen queries are expected, 2) choosing BEST when precision is crucial for repeated queries and the number of candidate n-grams is small, and 3) choosing LPMS for formatted datasets or when query literals are common in the dataset.
Based on this study, we also propose the following potential directions for future work in this area.
  1. (1)
    Unified solution. Our results demonstrate that while each method performs well on certain workloads, there is no solution that performs well across the board. Designing an algorithm that can combine the best aspects of each indexing method or prove the non-existence of such an algorithm would be beneficial.
  2. (2)
    Better indexing formats. Research in relation query processing has shown the benefits of using bit-based indexing formats (e.g. vector-based formats, BitWeaving (Li and Patel, 2013), etc.). It would be interesting to explore the benefits of those formats for regex indexing.
  3. (3)
    Indexing on-the-fly. All methods for regex indexing require an expensive preprocessing step. To the best of our knowledge, there are no existing indexing method that can perform indexing on-the-fly as the queries in the workload are executed and as the dataset and query distribution change dynamically. Adaptive query execution in this setting is another interesting direction for future work.
8. Conclusion
In this paper, we present an investigation of common regex indexing methods. Our comprehensive evaluation spans a diverse array of scenarios and datasets, establishing a benchmark for assessing the performance of these techniques. Through meticulous analysis, we have identified the inherent strengths and limitations of various methods. Based on our findings, we create a recommendation for practitioners on what method to choose based on the input characteristics. We also highlight the important open problems for regex indexing to spur further research in this area.

### v2

@@
-An Evaluation of N-Gram Selection Strategies for Regular Expression Indexing in Contemporary Text Analysis Tasks [Experiment, Analysis & Benchmark]
+An Evaluation of N-Gram Selection Strategies for Regular Expression Indexing in Contemporary Text Analysis Tasks. Extended Version
 Abstract.
@@
 Despite its widespread adoption, the scalability of this approach hinges on a critical problem: how to select the optimal set of n-grams to index, balancing trade-offs between index size, construction time, and query accuracy.
-Existing solutions, such as frequency-based selection (prioritizing common n-grams) (Cho and Rajagopalan, 2002), coverage-optimized methods (minimizing redundant matches) (Hore et al., 2004), and heuristic-driven strategies like LPMS (Tsang and Chawla, 2011), were proposed decades ago.
-For instance, the FREE (Cho and Rajagopalan, 2002) and BEST (Hore et al., 2004) methods, introduced in the early 2000s, focus on minimizing computational space overhead and index size for memory-constrained systems.
+Legacy methods like frequency-based selection (Cho and Rajagopalan, 2002), coverage-optimized strategies (Hore et al., 2004), and heuristics like LPMS (Tsang and Chawla, 2011) were introduced decades ago to reduce index size and memory usage (Cho and Rajagopalan, 2002; Hore et al., 2004).
 Despite being an active area of research, surprisingly, the performance of these methods on modern hardware is lacking for three reasons.
-The absence of a comprehensive, unbiased, and systematic understanding of these methods is due to three main reasons.
-First, existing methods have not been evaluated within the same experimental settings (such as the same hardware or using the same B+-Tree implementation), leading to incomplete comparisons.
-Second, prior work has only compared specific metrics and on a limited number of datasets.
-Third, the absence of a standardized and inclusive testing framework hinders the investigation and testing of methods in practical implementations.
-Consequently, the lack of a modern comparison of these past methods has led to the adoption of these legacy strategies in recent studies (Qiu et al., 2022; Chockchowwat et al., 2022) without re-evaluating their assumptions.
-These older methods may not be appropriate when dealing with larger datasets (now common in applications domains such as genenomics, IoT/sensor, and log analytics) as it stresses their n-gram selection methods.
-For example, as we will see later, BEST’s quadratic runtime complexity becomes prohibitive on even a small 10GB text corpus, forcing practitioners to downsample data or abandon indexing altogether.
-In fact, n-gram indexing used by Postgres (Christopher Kings-Lynne and Korotkov, [n.d.]) and GitLab code search (Peterse, 2016) abandon the n-gram selection strategies and build the index with all n-grams for only trigrams (i.e.
-$n=3$) or for several fixed values $n$ (for instance, indexing all bigrams and trigrams).
-In summary, there are three key limitations that motivate our work.
+The absence of a comprehensive, unbiased, and systematic understanding of these methods is due to three main reasons.: (1) inconsistent experimental setups hinder fair comparisons, (2) evaluations are narrow in scope and datasets, and (3) a lack of standardized testing framework hindering practical assessment.
+As a result, recent studies (Qiu et al., 2022; Chockchowwat et al., 2022) continue to adopt these outdated approaches without re-evaluating their assumptions.
+These methods struggle with today’s large datasets (e.g., genomics, IoT, logs), where selection strategies break down.
+For instance, BEST’s quadratic runtime becomes impractical on even 10GB of text, often forcing downsampling or skipping indexing.
+Modern systems like Postgres (Christopher Kings-Lynne and Korotkov, [n.d.]) and GitLab (Peterse, 2016) bypass selection entirely, indexing all trigrams or fixed n-grams.
+These limitations motivate our work.
   1. M.1
@@
 Prior evaluations focus on specific metrics (e.g., false-positive rate) or evaluate their method using only a single workload, failing to capture the diverse demands of real-world workloads.
-For example, genomic databases prioritize minimizing false positives due to a large number of repeated queries on the same dataset, while log analysis prioritizes rapid index rebuild overhead for large scale streaming data analysis.
 No study has systematically compared and contrasted the three key methods FREE, BEST, and LPMS across wide-ranging scenarios.
@@
 We modernized all three methods by: (1) redesigning FREE with parallelism to leverage modern multi-core CPUs, and (2) implementing LPMS with Gurobi, which inherently supports multi-threading in its LP solver.
-Additionally, we introduced an early stopping mechanism to all three methods to configure the maximum number of n-grams selected.
-We also developed an end-to-end benchmark framework that manages data loading, method selection, configuration settings, regex matching, and result reporting.
-This framework is modular, easily extensible, and publicly available at the link noted in the abstract.
+Our benchmarking framework is modular, easily extensible, and publicly available at the link noted in the abstract.
 All experiments were conducted on a multi-core machine with large memory, addressing (M.2).
@@
 Broad Workload Benchmark.
-To address the lack of comprehensive empirical comparisons on workloads with different characteristics, we evaluate the methods across five workloads.
+To address the lack of comprehensive empirical comparisons on workloads with different characteristics, we evaluate the methods across eight workloads.
 This includes using three legacy benchmarks from prior work and two real-world workloads in contemporary regex matching tasks.
@@
 Early works compared different n values and decided to index all trigrams.
+We refer to this method as Trigrams, and will use it for comparison in Section 6.
 Other works index n-grams with different n values for various workloads.
@@
 Similar ideas have also been used for indexing data to speed up regular path query evaluation (Li et al., 2001).
-The first known works discussing n-gram selection for regex indexing is FREE (Cho and Rajagopalan, 2002).
+The first known work discussing n-gram selection for regex indexing is FREE (Cho and Rajagopalan, 2002).
 The high-level idea of selecting a covering n-gram set was proposed in a study of Asian language indexing (Ogawa and Matsuda, 1998).
@@
 In this work, we compare the n-gram selection and indexing methods of FREE, BEST, and LPMS for in-memory workloads and indices.
+Several works have explored ML-based approaches for regex evaluation in security contexts.
+On resource-constrained edge devices, compressed on-device ML models are commonly used for high-throughput malicious packet detection.
+However, these methods rely on extensive preprocessing and assume access to high-quality training data—assumptions that don’t hold in log analytics, where analysis is routinely ad hoc.
+For a broader overview, see the survey by Xu et al.
+(Xu et al., 2016).
 3. Problem Definition
 In this section, we discuss the n-gram selection problem and provide formal definitions for some common terms.
-Definition 3.1.
-N-Gram For a given finite alphabet $\Sigma$, an n-gram $g$ is a sequence of $n$ characters $g=g_{1}g_{2}\cdots g_{n}$ where $g_{i}\in\Sigma$.
+Definition 3.1 (N-Gram).
+For a given finite alphabet $\Sigma$, an n-gram $g$ is a sequence of $n$ characters $g=g_{1}g_{2}\cdots g_{n}$ where $g_{i}\in\Sigma$.
 We use $|g|$ to denote the length of the n-gram $g$.
@@
 For a literal $l$, we will use $G(l)$ to denote the set of all possible substrings of $l$.
-The set of all possible n-grams of a regex $q$ is $G(r)=\bigcup_{\text{maximal }l\in q}G(l)$.
+The set of all possible n-grams of a regex $q$ is $G(q)=\bigcup_{\text{maximal }l\in q}G(l)$.
+We use $G$ to define the set of all candidate n-grams considered in an n-gram selection method.
+Note that $G$ may be different for different methods, and we will further discuss it in Section 4.
+$G_{i}$ to is used to denote subset of candidate n-grams that has length exactly $i$.
 Example 3.2.
@@
 The set of literal components in the example regex are: {<ahref=, ZZZ.pdf, >}.
-A workload $W=(Q,D)$ consists of the set of regex query $Q=\{q_{1},q_{2},\cdots\}$ and dataset $D=\{d_{1},d_{2},\cdots\},d_{i}\in\Sigma*$.
+A workload $W=(Q,D)$ consists of the set of regex queries $Q=\{q_{1},q_{2},\cdots\}$ and dataset $D=\{d_{1},d_{2},\cdots\},d_{i}\in\Sigma*$.
 The size of the dataset is defined as $|D|=\sum_{d_{i}\in D}|d_{i}|$.
@@
 Next, we define the support of an n-gram.
-Definition 3.3.
-Support The support $s$ of a literal $g$ in dataset $D$ is the number of elements in $D$ that contains $g$.
+Definition 3.3 (Support).
+The support $s$ of a literal $g$ in dataset $D$ is the number of elements in $D$ that contains $g$.
 Similarly, its support in query set $Q$ is the number of individual queries which contains $g$ as a literal.
@@
 Using the notion of support, we define the selectivity of an n-gram $g$ as follows.
-Definition 3.4.
-Selectivity The selectivity $c$ of an n-gram $g$ in string set $D$ is the fraction of individual strings in $D$ that contains $g$.
-$c_{g}=\frac{s_{D}(g)}{|D|}$
+Definition 3.4 (Selectivity).
+The selectivity $c$ of an n-gram $g$ in string set $D$ is the fraction of individual strings in $D$ that contains $g$, i.e., $c_{g}={s_{D}(g)}/{|D|}$.
 For n-gram selection methods, selectivity is an important parameter when deciding if an n-gram is selected for indexing or not.
@@
 Table 1. Selection methods summary.
-Method N-Gram Source Selection Criteria Selectivity Threshold ($c$) N-Gram Constraint Index Structure Index Size Constraint
-FREE (Cho and Rajagopalan, 2002) dataset prefix-free, selectivity 0.1 $2\leq n\leq 10$ Inverted Index -
-BEST (Hore et al., 2004) queries & dataset Utility(g) = benefit($g$)/cost($g$) $0.05\leq c\leq 0.1$ B+-Tree Index User defined
-LPMS (Tsang and Chawla, 2011) queries & dataset prefix-free and utility optimized - - Inverted Index -
+Method Source Selection Criteria Selectivity Threshold ($c$) N-Gram Constraint Time Complexity Space Complexity
+FREE (Cho and Rajagopalan, 2002) $D$ Prefix-free Selectivity 0.1 $[2,10]$ $O(k\cdot|D|)$ $O(|D|)$
+BEST (Hore et al., 2004) $Q,D$ Utility $[0.05,0.1]$ - $O(|D|+|Q|)$ $O(|I|\cdot|G(Q)|\cdot|D|\cdot|Q|)$
+LPMS (Tsang and Chawla, 2011) $Q,D$ Prefix-free Utility optimized - - $O(\sum_{i\in[10]}\{|G_{i}(Q)|^{2.5}$ $\quad+|G_{i}(Q)|\cdot(|D|+|Q|)\})$ $O(\max_{i\in[10]}|G_{i}(Q)|\cdot(|D|+|Q|)$
+VGGraph (Qiu et al., 2025) $Q,D$ Combines FREE and BEST - $2,\inf)$ $O(n\cdot|D|+|Q|\cdot|G(Q)|\cdot n)$ $O(|D|+|G(Q)|)$
 4.1. FREE
@@
 Note that FREE does not use the query workload for n-gram selection.
-The candidate set $G(W)$ of n-grams is $G(W)=\bigcup_{d\in D}G(d)$.
-For the candidates $G(W)$, FREE decides whether a n-gram will be selected by Usefulness.
-Definition 4.1 (FREE).
-Usefulness By setting a fixed selectivity threshold $\mathbf{c}$, n-gram $g$ is deemed useful if its selectivity is less than the selectivity threshold, $c_{g}<\mathbf{c}$.
-FREE makes two important assumptions when selecting n-grams for indexing.
+The candidate n-gram set $G$ is $G(D)=\bigcup_{d\in D}G(d)$.
+FREE decides whether an n-gram will be selected by the Usefulness criteria as defined below.
+Definition 4.1 (FREE Usefulness).
+By setting a fixed selectivity threshold $\mathbf{c}$, n-gram $g$ is deemed useful if its selectivity is less than the selectivity threshold, $c_{g}<\mathbf{c}$.
+FREE makes two important assumptions for n-grams selection.
 Assumption FREE-1 0.
@@
 FREE selects only n-grams that are useful.
-Since usefulness selection criteria is essentially a selectivity upper bound, if we know that a n-gram $g$ is useful, then any longer n-gram having $g$ as a substring is also useful.
-For example, if n-gram pdf is useful with a selectivity $c_{g}$, then n-grams such as $g^{\prime}=$pdf and $g^{\prime\prime}=$ pdfg will have selectivity $0\leq c(g^{\prime})<\mathbf{c}$ and $0\leq c(g^{\prime\prime})<\mathbf{c}$, and therefore also useful.
+Since usefulness selection criteria is essentially a selectivity upper bound, if we know that an n-gram $g$ is useful, then any longer n-gram having $g$ as a substring is also useful.
+For example, if n-gram pdf is useful with a selectivity $c_{g}$, then n-grams such as $g^{\prime}=$ .pdf and $g^{\prime\prime}=$ pdfg will have selectivity $c(g^{\prime})<\mathbf{c}$ and $c(g^{\prime\prime})<\mathbf{c}$, therefore also useful.
 Property FREE: Usefulness 0.
-For an n-gram $g=g_{1}g_{2}\cdots g_{n_{1}}$ of size $n_{1}$ having selectivity $c_{g}$, any other longer n-grams of size $n_{2}$
+For a useful n-gram $g=g_{1}g_{2}\cdots g_{n_{1}}$ of size $n_{1}$ having selectivity $c_{g}$, any other longer n-grams of size $n_{2}$
 $g^{\prime}=p_{1}\cdots p_{m}g_{1}g_{2}\cdots g_{n_{1}}s_{1}\cdots s_{n_{2}-n_{1}-m}$
@@
 Essentially, it generates all candidate n-grams in increasing size order iteratively.
-For iteration $i$, it generates the candidate set of n-grams of length $i$ by extending all useless n-grams of length $i-1$ by one character, inserting the useful n-grams in the candidate set into the index, and using the useless ones in the candidate set for iteration $i+1$.
+For iteration $i$, it generates the candidate set of n-grams of length $i$ by extending all useless n-grams of length i1i-1 by one character, inserting the useful n-grams in the candidate set into the index, and using the useless ones in the candidate set for iteration $i+1$.
 This way, it is not necessary to generate all possible n-grams in each iteration.
@@
 In the original paper, the authors uses $c=0.1$ and $2\leq n\leq 10$ in their experiments.
-The accompanied index data structure used is inverted index with n-grams as keys and posting lists as values.
-(a) Literal based query plan before checking index keys.
-(b) Final query plan after index scan.
-Figure 1. Example index search plan tree for regex <ahref=("|’).*ZZZ\. pdf("|’)> built by FREE, where the n-grams ", , Z, and pdf are indexed.
-4.1.2. Regex Compilation
-In order to also handle regex queries literals in alternative strings, FREE include a simple regex query plan compiler that compiles the overall index lookup to a plan tree with only logical AND and OR operators.
-Specifically, for a given regex, the compiler identifies literals on the primary level and literals, and then build a tree with the literals.
-FREE then checks if each of the literals or parts of the literals are indexed in the index, removing subtrees with no literals indexed, and substituting literal nodes with index keys, which are n-grams in the index.
-We provide an example index search query plan in Figure 1 for the example regex, <ahref=("|’).*ZZZ\.
-pdf("|’)>, assuming that only ", , Z, and pdf are indexed.
-The literal tree shown in Figure 1(a) is built sorely by extracting all literals from the regex.
-Referring to the index, we notice that <ahref= and > are not indexed, and two n-grams in the literal ZZZ.pdf, Z and pdf are indexed.
-We simplify the plan to get Figure 1(b).
-The evaluation of the index search plan tree then is a depth-first evaluation, where AND is a set intersection operation of two posting lists and OR is a set union operation of two posting lists.
+It also defined a simple query plan considering the index keys.
+We include a description of the query plan generation in Appendix A of the full paper.
+We implemented the original version of FREE with the query plan and an accompanied matcher in the codebase, but for a fair comparison of the n-gram selection methods, we do not use it for the experiments.
+The index data structure used is inverted index with n-grams as keys and posting lists as values.
 4.1.2. Complexity Analysis
-Let us denote the candidate set of each step $i$ as $M_{i}$.
-In the initial step, FREE scans the entire dataset to find all unique uni-grams and calculates their selectivities.
-The space needed is a temporary hashmap with the key being all unique unigrams of size $|M_{1}|=|\Sigma|$.
-The time complexity for the base step is $O(|D|+|\Sigma|)$.
-After scanning through the dataset, a pass is made on all unique unigrams to calculate their selectivity and move the useful unigrams from the candidate set into the index.
-In the next iteration, the remaining useless uni-grams in the candidate set of the previous step are extended by one character to construct a candidate set of bi-grams, $M_{2}$.
-The runtime space overhead of this iteration is $O(|M_{2}|)$, and the compute time overhead is $O(|D|+|M_{2}|)$.
-For each iteration $i$, we have the runtime space overhead $O(|M_{i}|)$ and compute overhead $O(|D|+|M_{i}|)$.
-Since each step $i$ selects all useful n-grams of size $i$, and the temporary hashmaps of previous iterations are never revisited after candidate set $i$ is generated, for a FREE n-gram selection with $n\leq k$ for some integer $k$, we have the runtime space overhead $O(|M_{k}|)$ and compute overhead $O(k\cdot(|D|+|M_{k}|))$ for each iteration.
-As the number of n-grams of any size on a dataset is at most the number of characters in the dataset, we have the overall runtime space overhead $O(|D|)$ and the compute overhead as $O(2\cdot|D|)$ for all iterations.
+Let the candidate set at step $i$ be $M_{i}$.
+Initially, FREE scans the dataset to identify all unique unigrams and compute their selectivities, using a temporary hashmap of size $|M_{1}|=|\Sigma|$.
+This base step has time complexity $O(|D|+|\Sigma|)$.
+After filtering for useful unigrams, the remaining ones are extended to form bigrams $M_{2}$, with space and compute overheads of $O(|M_{2}|)$ and $O(|D|+|M_{2}|)$, respectively.
+At each step $i$, the runtime space is $O(|M_{i}|)$ and compute time is $O(|D|+|M_{i}|)$.
+Since previous hashmaps are discarded, selecting n-grams up to size $k$ results in total space $O(|M_{k}|)$ and compute time $O(k\cdot(|D|+|M_{k}|))$.
+Given that $\sum_{i=1}^{k}|M_{i}|\leq|D|$, the overall space is $O(|D|)$ and compute time is $O(k\cdot|D|)$.
+In practice, $k$ is usually small and can be treated as a constant.
 4.2. BEST
@@
 Formally, we have the following construction.
-Definition 4.3 (BEST).
-Cover For a workload $W=(Q,D)$, the cover of an n-gram $g$ is the set of data records $d\in D$ and query $q\in Q$ pairs defined as follows:
-$cover(g)=\{(q,d)\quad|\quad g\in q\land g\notin d,\quad\forall(q,d)\in Q\times D\}$
-For a set of n-grams $G$, the cover of the n-gram set is the union of the cover of each n-gram in the set.
-$cover(G)=\bigcup_{g\in G}cover(g)$
-The n-gram selection problem is then transformed to a budgeted maximum set cover problem (Khuller et al., 1999) that aims to find the minimum set $G$ that achieves maximum cover.
-Formally, BEST define the value of indexing an n-gram with respective to the entire workload by its benefit.
-Definition 4.4 (BEST).
-Benefit For a workload $W=(Q,D)$ and an n-gram set $I$, the benefit of an n-gram $g\notin I$ is the number of additional query-data pairs covered by $g$ that is not already covered by $I$.
+Definition 4.3 (BEST Cover).
+For a workload $W=(Q,D)$, the cover of an n-gram $g$ is the set of data records $d\in D$ and query $q\in Q$ pairs defined as follows:
+$cover(g)=\{(q,d)\in Q\times D\mid g\in q\land g\notin d\}$
+For a set of n-grams $P$, the cover of the n-gram set is the union of the cover of each n-gram in the set.
+$cover(P)=\bigcup_{g\in P}cover(g)$
+The n-gram selection problem is then transformed to a budgeted maximum set cover problem (Khuller et al., 1999) that aims to find the set that maximize the cover over all sets $P$ that satisfy the budget constraint Formally, BEST define the value of indexing an n-gram with respective to the entire workload by its benefit.
+Definition 4.4 (BEST Benefit).
+For a workload $W=(Q,D)$ and an n-gram set $I$, the benefit of an n-gram $g\notin I$ is the number of additional query-data pairs covered by $g$ that is not already covered by $I$.
 $benefit(g,I)=|cover(I\cup\{g\})-cover(I)|$
-Definition 4.5 (BEST).
-Cost For a workload $W$ consists of dataset $D$ and query set $Q$, the cost to index an n-gram $g$ is the storage overhead of $g$.
+To progressively calculate the benefit of each n-gram $g$ of $G(W)=G(D)\bigcup G(Q)$, we need two matrices to store the existence of $g$ in each data record $d$ and each query $q$.
+Specifically, the two matrices are of size $|Q|\cdot|G(W)|$ and $|D|\cdot|G(D)|$ respectively.
+Definition 4.5 (BEST Cost).
+For a workload $W$ consists of dataset $D$ and query set $Q$, the cost to index an n-gram $g$ is the storage overhead of $g$.
 The cost is dependent of the index structure.
@@
 For inverted index, the cost of $g$ is the size of its list, which is of size $s_{D}(g)$.
-Definition 4.6 (BEST).
-Utility For a workload $W$ consists of dataset $D$ and query set $Q$ and an existing n-gram set $I$ on $W$, the utility of indexing an additional n-gram $g$ is the ratio of its benefit over cost.
-$utility(g)=\frac{benefit(g,I)}{cost(g)}$
-BEST selects n-grams based on their utility as defined above.
-The brute-force method would be to iteratively select the n-gram with the highest utility, $g_{max}$, based on the workload $W$ and the set of already-selected n-grams $I$, among all possible n-grams with a positive benefit.
+Definition 4.6 (BEST Utility).
+For a workload $W$ consists of dataset $D$ and query set $Q$ and an existing n-gram set $I$ on $W$, the utility of indexing an additional n-gram $g$ is the ratio of its benefit over cost, i.e., utility(g)=benefit(g,I)/cost(g)utility(g)={benefit(g,I)}/{cost(g)}
+BEST selects n-grams based on their utility.
+The brute-force method would be to iteratively select the n-gram with the highest utility, $g_{max}$, based on the workload $W$ and the set of selected n-grams $I$, among all possible n-grams with a positive benefit.
 Assumption BEST-3 0.
@@
 The average number of characters in all literals of a regex is much smaller than that of data records. The number of regex queries is much smaller than number of data records in the same workload.
-Therefore, according to the sparsity assumption Assumption BEST-3, BEST choose to use adjacency lists: Q-G-list and G-D-list rather than two matrices of sizes $|Q|\cdot|G|$ and $|G|\cdot|D|$ to store the existence of n-grams in the workload to reduce space usage.
+Therefore, according to the sparsity assumption Assumption BEST-3 and the definition of cover, instead of considering all n-grams in the query set and dataset, we can trim away the n-grams that only exists in the dataset.
+Therefore, the candidate n-gram set $G$ of BEST is $G(Q)$.
+BEST choose to use adjacency lists: $Q$-$G$-list and $G$-$D$-list rather than two matrices of sizes $|Q|\cdot|G(W)|$ and $|G(W)|\cdot|D|$ to store the existence of candidate n-grams to reduce space usage.
 The candidate n-grams, queries, and data records are each assigned an index number.
@@
 Besides, the search space is the product of number of possible query-data pairs, $|Q|\cdot|D|$ and the number of possible n-grams, $|G|$.
-By Assumption Assumption BEST-4, we assume $G=G(Q)$.
-When the query size and the dataset size is large, the number of candidate n-grams can be prohibitively large to select the n-grams using brute-force method.
+When the query size and the dataset size is large, the number of candidate query-data pairs can be prohibitively large to select the n-grams using brute-force method.
 BEST employs several techniques to get an approximate result.
@@
 4.2.3. Complexity Analysis
-First, the workload reduction technique is applied, reducing the query set $Q$ to a reduced set $|Q^{\prime}|=\frac{|Q|}{t}$ for some $t\geq 1$ and the candidate n-gram set being $G^{\prime}$.
-It is reasonable to assume that $|G^{\prime}|\simeq\frac{|G|}{t}$.
-Let $\mu$ be the average number of n-grams in a query and $T$ be the maximum number of iterations of the k-median algorithm.
-Then, the workload reduction time is $O(|Q|\cdot|Q^{\prime}|\cdot\mu\cdot T)$.
-We use a suffix tree to efficiently enumerate all possible n-grams in the queries, which will take $O(|G^{\prime}|)$ time to build, enumerate, and insert them to a hash-map.
-A similar clustering is then carried out for creating sub-problems for parallelism.
-Empirically, according to the original BEST paper, dividing the problem to sub-problems reduces both time and space complexity by approximately $5$ even if we run it on a single thread.
-The use of the two adjacency lists Q-G-list and G-D-list rather than two matrices reduces the space overhead for utility calculation from $O(|G|\cdot(|D|+|Q|))$ to $O(\mu\cdot(|D|+|Q|))$.
+First, the workload reduction step shrinks the query set from $|Q|$ to $|Q^{\prime}|=\frac{|Q|}{t}$, with the candidate n-gram set $G^{\prime}$ also reduced proportionally.
+Assuming average query length $\mu$ and $T$ iterations of k-median, this step takes $O(|Q|\cdot|Q^{\prime}|\cdot\mu\cdot T)$ time.
+A suffix tree is used to enumerate all n-grams in $O(|G^{\prime}|)$ time.
+Next, clustering enables parallelism and, per the original BEST paper, reduces time and space by $\sim$5 $\times$ even on a single thread.
+Using adjacency lists (Q-G-list and G-D-list) instead of matrices cuts space from $O(|G|\cdot(|D|+|Q|))$ to $O(\mu\cdot(|D|+|Q|))$, which is small since $\mu\ll|G|$.
+However, building these lists still takes $O(|G|\cdot(|D|+|Q|))$ time.
 By Assumption Assumption BEST-4, $\mu\ll|G|$, and therefore, the memory overhead during computation is small.
@@
 For each pair $(q,d)\in(Q^{\prime}\times D)$, it checks if the pair covered by each candidate n-gram $g$.
-The time complexity of each iteration is $O(|G^{\prime}|\cdot|Q^{\prime}|\cdot|D|)$.
-The overall time complexity for BEST including workload reduction and clustering can be expressed as $O(|Q|\cdot|Q^{\prime}|\cdot\mu\cdot T+|G|\cdot(|D|+|Q|)+\frac{|I|\cdot|G^{\prime}|\cdot|Q^{\prime}|\cdot|D|}{5})$, which simplifies to $O(\frac{|I|}{5t^{2}}\cdot|G|\cdot|Q|\cdot|D|)$.
+In each iteration, BEST evaluates all remaining n-grams against $(q,d)\in Q^{\prime}\times D$, costing $O(|I|\cdot|G(Q)|\cdot|Q|\cdot|D|)$.
 4.3. LPMS
@@
 Formally,
-Definition 4.7 (LPMS).
-Coverage The coverage of an n-gram $g$ is defined as the ratio of the support of $g$ in the dataset $D$ and the support of $g$ in the query set $Q$, normalized by the n-gram length.
-$cv(g)=\frac{s_{D}(g)}{|g|\cdot s_{Q}(g)}$
+Definition 4.7 (LPMS Coverage).
+The coverage of an n-gram $g$ is defined as the ratio of the support of $g$ in the dataset $D$ and the support of $g$ in the query set $Q$, normalized by the n-gram length: $cv(g)={s_{D}(g)}/{|g|\cdot s_{Q}(g)}$
 Similar to Assumption Assumption BEST-4, we assume $G=G(Q)$.
@@
 minimize $\sum_{g\in G}cv(g)x_{g}$
-subject to $Ax\geq b$
-$0\leq x_{g}\leq 1\quad\forall g\in G$
+subject to $Ax\geq b;0\leq x_{g}\leq 1\quad\forall g\in G$
 4.3.2. Complexity Analysis
@@
 In practice, the size of the index key typically does not exceed 10, a number also used as the upper bound for n-gram size in FREE.
-Therefore, we can consider the small number of iterations as a constant, making the overall computational complexity of LPMS $O(\sum_{i\in[10]}|G_{i}|^{2.5}+|G_{i}|\cdot(|D|+|Q|))$.
+Therefore, we can consider the small number of iterations as a constant, making the overall computational complexity of LPMS $O(\sum_{i\in[10]}\{|G_{i}|^{2.5}+|G_{i}|\cdot(|D|+|Q|)\})$.
+4.4. VGGraph
+The n-gram selection strategy from VGGraph (Qiu et al., 2025) combines the insights from FREE and BEST to select the n-grams.
+4.4.1. Selection Strategy
+The idea is to select the most cost-effective n-gram from a set $S$ to cover characters in an uncovered set $E$.
+Each n-gram is evaluated based on a cost-efficiency ratio (a notion similar to cover from Definition 4.3), where the cost is size of the inverted index for the n-gram and the efficiency is the number of characters it helps cover in $E$.
+VGGraph approximate a perfect set cover with efficiency by adopting FREE’s idea of Usefulness, and only extend the useful n-grams.
+By maintaining the uncovered set $E$ that is continuously updated as the n-grams are picked, VGGraph ensures that prefix-free n-grams are selected preferentially.
+4.4.2. Complexity Analysis
+VGGraph’s index key selection consists of two steps: (1) recursive extension of grams exceeding the frequency threshold $c$, and (2) greedy n-gram selection via the set-cover approximation.
+Given a dataset $D$, the complexity of recursively extending grams is $O(|D|\cdot q_{\min})$, as each character position is indexed exactly once and each gram is extended at most once.
+The greedy set-cover step selects, for each regex, a minimal subset of grams from the candidate set $G(Q)$.
+For a regex with literal size $n$, the greedy selection has a complexity of $O(|G(Q)|\cdot n)$ per query.
+The total runtime complexity for key selection, combining both steps, is thus $O(|D|\cdot q_{\min}+|Q|\cdot|G(Q)|\cdot n)$.
+The corresponding runtime space complexity during key selection is $O(|D|+|G(Q)|)$, dominated by storing candidate grams and their counts.
 5. Experiment Setup
 Due to the absence of source code from the papers, we implemented the three n-gram selection methods ourselves.
-For each method, we also developed their corresponding query parsers (if any) and index structures.
+For a fair comparison of the n-gram selection methods, we use the same index structure (inverted index) and same basic query matcher for all experiments.
+We include the implementations of other index structures and query plans per the original papers in our codebase for reference.
 5.1. Benchmark Framework
@@
 The end-to-end process follows a seven-step sequence, as illustrated in the framework diagram Figure 1.
-First, the user specifies inputs (dataset, regex queries, method name, and configurations), which are then processed by the framework.
-The selected n-gram strategy processes the workload, identifying optimal n-grams for index construction.
-These n-grams are used to build an index (e.g., inverted index or B+-tree).
-The index search plan compiles by examining the regex patterns and extracting the n-grams in their literals to create plans for index lookup, filtering candidate data points.
-The regex engine then verifies these candidate data points, discarding false positives, and generating metrics to quantify performance.
-Finally, the results are aggregated and exported.
-The framework operates in a pipeline architecture divided into three phases: input processing, index construction, and regex evaluation.
-The components of each phase are grouped within dashed squares in Figure 2.
+t begins with user-specified inputs (dataset, regex queries, method, and configurations), which are processed by the framework.
+The selected n-gram strategy analyzes the workload to identify optimal n-grams for index construction.
+These n-grams are then used to build an index (e.g., inverted index or B+-tree).
+Next, the index search plan extracts n-grams from regex literals to guide candidate filtering.
+The regex engine verifies these candidates, discards false positives, and computes performance metrics.
+Finally, results are aggregated and exported.
+The pipeline is organized into three phases—input processing, index construction, and regex evaluation—with components grouped accordingly in Figure 1.
 Input Processing.
@@
 In the index construction phase, one of the three n-gram selection strategies is applied to the workload.
-The selected n-grams are used to build an index, such as an inverted index or a B+-tree.
+The selected n-grams are used to build an index.
 This phase leverages multi-threaded execution if specified by the user.
 After the index is built, the index search plan is compiled if necessary.
-Each strategy has its own corresponding type of index structure and index search plan, as shown in the blue dashed bounding box in Figure 2.
-FREE and LPMS use inverted indexes, while BEST uses a B+-tree for the index structure.
-During index lookup, all n-grams in both the set of regex query literals and index keys are extracted.
-Only FREE requires an additional intermediate step to generate an index search tree.
+An inverted index with data string IDs is built with the selected n-grams.
+In the simple query plan, posting lists corresponding to the set of n-grams that exists in the regex query literals are unioned together later for full regex evaluation.
 Regex Evaluation and Result Output.
@@
 We use $\bar{|d|}$ to denote average data record size (in bytes) and $\bar{TP}$ to denote the average number of actual matches of all queries.
-Synthetic has two query sets – the first set of 500 queries that is used to build the index, and a second set of 100 queries that is used to test the impact of queries that are not indexed.
-Workload $|Q|$ $|D|$ ||$\sum$|| ∑ | $\bar{|d|}$ $\bar{TP}$ Dataset Size
+Robust has two query sets – the first set of 500 queries that is used to build the index, and a second set of 100 queries that is used to test the impact of queries that are not indexed.
+Workload $|Q|$ $|D|$ $|\sum|$ $\bar{|d|}$ $\bar{TP}$ Dataset Size
 Webpages 10 695,565 255 68,650 86,524 47 GB
@@
 SQL-Srvr 132 101,876,733 114 139 50,356 14 GB
-Synthetic 500/100 5,000 16 32 628 163 KB
+Enron 107 517,401 120 2719 54,242 2.6 GB
+ADX 18 890,623,051 118 101 4,653,057 100 GB
+Synthetic vary 50,000 16 200 vary 9.6 MB
 5.2. Workloads
@@
 Among the workloads, three (Webpages, DBLP, and Prosite) are those used by the original FREE, BEST, and LPMS papers to evaluate their methods.
-The other two workloads (US-Accand SQL-Srvr) are more recent and feature larger datasets.
+The other four workloads (US-Acc, SQL-Srvr, Enron, ADX) are more recent and feature larger datasets.
 Webpages.
@@
 SQL-Srvr. This is a production workload consisting of in total 101,876,733 log messages generated by Microsoft SQL Server and 132 regex queries used for data analysis tasks on this dataset.
+Enron.
+This is a production workload consists of more than 50 thousand of real emails and 107 regex queries.
+The dataset is the May 7, 2015 Version of Enron dataset (Cohen, 2015), and the queries are from Avatar search engine (Jayram et al., 2006) collected by Chen et al.
+(Chen, 2012).
+ADX. This is a real production workload consists of 890 millions of logs from Azure Data Explorer (ADX) and 18 regexes for an log analysis task on this dataset by an data scientist.
 Synthetic.
-To test the robustness of the n-gram selection methods in handling unseen queries, we build off the synthetic dataset in the original LPMS (Tsang and Chawla, 2011) paper.
-The dataset consists of 5000 strings each constructed with alphabet ‘A-P’.
-The string size follows a geometric distribution with $p=1/32$.
-The index building query set is generated by randomly selecting 10% from the dataset, and a random slice is used to create a regex lit1regexlit2.
-lit1 has 1-5 characters, lit2 has 0-5 characters, and regex matches any $m$ characters where $1\leq m\leq 50$.
-During the query phase, we generate regexes of a similar format from 2% of the data records.
+This workload is based on the method described in the paper proposing LPMS (Tsang and Chawla, 2011).
+The workload consists of 50,000 data records with a fixed alphabet size of 16.
+We varied the query set size across $[500,1000,2000,5000]$ and explored different levels of query selectivity ranging from 0.001 to 0.5.
 5.3. Metrics
@@
 $Prec_{W}=\frac{\sum_{q\in Q}\#TP_{q}}{\sum_{q\in Q}\#TP_{q}+\sum_{q\in Q}\#FP_{q}}$
-Workload Matching Time ($T_{Q}$). We measure the time to run the workload, denoted by $T_{Q}$, to demonstrate the runtime gains provided by the index.
+Query Time ($T_{Q}$). We measure the time to run the workload, denoted by $T_{Q}$, to demonstrate the runtime gains provided by the index.
 Runtime Space Usage ($S_{Q}$). This metric is the peak space used when running a experiment.
-Index Size ($S_{I}$). We measure the size of the index constructed, which is the total size of the n-gram keys, the posting lists, and necessary index data structure (such as the B+-tree).
+Index Size ($S_{I}$). We measure the size of the index constructed, which is the total size of the n-gram keys, the posting lists, and necessary index data structure (such as the inverted index).
 5.4. Hardware and Implementation Details
-We implemented FREE and its query parser, integrating it with an inverted index.
-For LPMS, we also used the same inverted index structure to ensure consistency.
-We implemented BEST with a B+-tree as the supporting index data structure.
+We implemented n-gram selection methods FREE, BEST, LPMS, VGGraph, and Trigrams with the index construction and matching methods mentioned in the benchmarking framework.
+We implemented FREE’ query parser and BEST with a B+-tree in the codebase as well.
 We used Gurobi Optimizer (Gurobi Optimization, LLC, 2024) as the linear program solver.
@@
 Google’s RE2 (release version 2022-06-01) (Google, [n.d.]) as the regex engine for regex evaluation.
-Each of the three methods has different configurable parameters.
-For BEST and FREE, we vary the selectivity threshold $c$ by picking values between 0.01 and 0.7.
-For FREE, we also vary the upper bound of n-gram length, $\max_{n}$, among values: 2, 4, 6, 8, and 10.
-Another parameter we vary in FREE is whether the selected n-gram set is prefix minimal or pre-suf minimal.
-We set a hard threshold of 3 hours for all experimental runs.
-If BEST fails to finish within this time limit, we reduce the workload by a factor of 0.1%, 0.2%, 0.5%, and 0.85%.
-LPMS allows us to configure whether to use deterministic relaxation (LPMS-D) or random relaxation (LPMS-R) of the linear programs.
-Similarly, we set a maximum number of n-grams for LPMS in case a workload takes too long to run.
-For all configurations of all methods, we use 16 threads.
-To fairly assess the methods, we compare the resulting indexes with a similar number of n-grams selected.
-Since FREE and LPMS do not natively support dynamic constraints on the index size, we implement an early-stopping mechanism for all three methods.
+Each method has its own configurable parameters.
+For BEST and FREE, we vary the selectivity threshold $c$ between 0.01 and 0.7.
+FREE also varies the maximum n-gram length ($\max_{n}\in\{2,4,6,8,10\}$) and whether the selected n-grams are prefix-minimal or pre-suf-minimal.
+All experiments are capped at 3 hours; if BEST exceeds this, we reduce the workload to 0.1–0.85%.
+LPMS supports both deterministic (LPMS-D) and randomized (LPMS-R) relaxations.
+We also cap the number of n-grams for LPMS if runtime is excessive.
+All methods run with 16 threads.
+To ensure fair comparison, we match index sizes across methods.
+Since FREE and LPMS don’t natively support key limits, we implement early stopping via a max-key parameter for all three methods.
 By providing an optional parameter of max number of keys, the n-gram selection methods will stop once the limit is reached.
-We choose the values of max number of keys on a case-by-case basis for each workload, as specified in Section 6.1.
-6. Experiments
-In this section, we run the aforementioned workloads on the three indexes with varying parameters. In particular, we aim to answer the following questions:
-  1. Q.1
-    How does each method perform on different datasets?
-  2. Q.2
-    How do the characteristics of each workload affect the usefulness of n-gram selection methods?
-6.1. Query Performance
-We determine the effectiveness of the index by evaluating the filtering effectiveness of the selected set of n-grams.
-The higher the precision of an index, the more effective the set of n-grams selected.
 Each configuration of a selection method generates a specific set of n-grams of varying sizes.
 To ensure a fair comparison, we compare selected sets of similar sizes, as indexing all possible n-grams would yield the highest precision for the workload.
+Since each methods has different configurations, and those configuration values interfere with each other in a non-linear way, isolating one configuration and study its behavior as value changes is not meaningful.
 For each n-gram selection method under a specific key number constraint $K$ $s.t.$ $|I|\leq K$, we select the configuration with the highest precision.
-Note that throughout the section, we report the numbers for space usage and query performance for the configuration that achieves the best possible precision given a constraint on $K$.
+We choose the values of max number of keys, $K$, on a case-by-case basis for each workload, as specified in Section 6.
+6. Experiments
+In this section, we run the aforementioned workloads on the indexes build from different n-gram selection methods with varying parameters. In particular, we aim to answer the following questions:
+  1. Q.1
+    What is the usefulness of each n-gram selection method? We evaluate this through precision and $T_{Q}$.
+  2. Q.2
+    What is the index construction overhead for each method? We evaluate this through $T_{I}$, $S_{Q}$, and $S_{I}$.
+  3. Q.3
+    How do the characteristics of each workload affect the usefulness of n-gram selection methods?
+We answer question Q.1 in Section 6.1 and question Q.2 in Section 6.2.
+To answer question Q.3, we discuss the impact of characteristics of different real-world workloads on index construction overhead in Section 6.2 and on index effectiveness in Section 6.1.
+We also use present a more thorough ablation study on impact of workload parameters such as the query set size $|Q|$, dataset size $|D|$, , query selectivity, etc.
+in Section 6.3.
+(a) Precision
+(b) Query Time $T_{Q}$ (s)
+Figure 2. DBLP.
+6.1. Query Performance
+To address question Q.1, we compare the performance gains from using indices built with n-grams selected by the three methods, focusing on filtering precision.
+Higher precision indicates a more effective n-gram set.
+We also measure the regex query runtime $T_{Q}$, which reflects precision since the index structure and matching strategy remain constant.
+Generally, indexing more n-grams improves precision.
+However, each method behaves differently.
+BEST often starts with the highest precision when the key budget $K$ is small, but its precision gains plateau due to time constraints (e.g., 3-hour index build).
+In contrast, FREE begins with lower precision—especially for small query sets—but improves significantly as $K$ increases, often achieving the highest precision overall.
+Precision of VGGraph is low when number of keys is small, but quickly rises as number of keys gets larger.
+LPMS has similar precision as BEST when the numbers of key is small and usually grows insignificantly as the number of n-grams selected increases.
+Trigrams typically has very low precision when the number of index keys fixed is low.
+We now investigate the methods’ precision for different workloads with varying characteristics.
 6.1.1. Workload DBLP
-Taking the DBLP workload as an example, we run BEST, FREE, and LPMS with varying parameters.
-After obtaining the results, we compare the performance of each n-gram selection method based on the upper bounds of the number of n-grams (i.e.
-index entries in $I$).
-Specifically, for this workload, we examine indices with a number of keys is at most $5000$.
-We summarize the results in Section 6.1.1, where the metric columns are variables introduced in Section 5.3.
-Table 3. Index cost and query performance on DBLP.
-$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
-150 BEST 19.7 533 2.194 33.433 0.235
-FREE 25.3 2 0.171 37.299 0.051
-LPMS 19.8 3 0.155 34.028 0.048
-500 BEST 15.2 8762 2.670 54.306 0.242
-FREE 20.6 2 0.242 66.774 0.165
-LPMS 17.7 6 0.158 40.657 0.070
-1000 BEST 15.2 8762 2.670 54.306 0.242
-FREE 18.7 1 0.287 75.354 0.219
-LPMS 17.0 808 1.309 41.499 0.071
-2000 BEST 15.2 8762 2.670 54.306 0.242
-FREE 17.4 1 0.264 68.442 0.321
-LPMS 17.0 808 1.309 41.499 0.071
-3000 BEST 15.2 8762 2.670 54.306 0.242
-FREE 16.9 2 0.304 83.069 0.524
-LPMS 17.0 808 1.309 41.499 0.071
-4000 BEST 15.2 8762 2.670 54.306 0.242
-FREE 16.9 2 0.304 83.069 0.524
-LPMS 17.0 808 1.309 41.499 0.071
-5000 BEST 15.2 8762 2.670 54.306 0.242
-FREE 16.6 2 0.316 85.266 0.535
-LPMS 17.0 808 1.309 41.499 0.071
-[…]
+We evaluate BEST, FREE, LPMS, and VGGraph across various input configurations, and applying the same key limit to Trigrams.
+Performance is compared based on index size, focusing on configurations with at most 5000 n-grams.
+As shown in Figure 2(a), BEST($c=0.5$) achieves a precision of 0.235 for $K\leq 150$, outperforming FREE($\max_{n}=2$, $c=0.5$) and LPMS-D, which yield much lower precision (0.051 and 0.048, respectively) at a similar index size.
+At $K\leq 500$, the 363 n-grams selected by BEST($c=0.5$) achieves significantly better precision compared to LPMS-D.
+With similar index construction overhead, FREE($\max_{n}=2$,$c=0.5$) achieves higher precision than LPMS-D.
+LPMS tends to select infrequent n-grams, which may not benefit overall workload filtering.
+Although BEST improves slightly with more keys (e.g., +0.007 precision and +4.5s query performance from $K=150$ to $K=500$), gains are modest.
+Meanwhile, VGGraph($\max_{n}=2$,$c=0.5$) builds an index with 215 n-grams but with a low precision of 0.035.
 When relaxing the index size constraint to 1000 n-grams, FREE ($\max_{n}=4$,$c=0.5$) generated an index of size 810 that achieved a precision of 0.219, slightly lower than BEST($c=0.5$).
-However, this configuration has a construction time of around 1 second, which is several orders of magnitude lower than both BEST and LPMS-D.
-FREE also consumes significantly less memory during computation.
-It uses 0.287 GB, whereas LPMS requires 1.3 GB and BEST 2.67 GB of memory to store the data and perform computation.
-For LPMS, with more n-grams selected, its second configuration achieves 10 times less index construction time and higher precision compared to using a smaller number of keys, resulting in a workload evaluation performance of 17.1 seconds, which is similar to the other methods.
-Although LPMS selects 929 keys to index, which is larger than the 810-key inverted index by FREE, it still has a smaller index size.
-This again demonstrates that LPMS tends to select n-grams with low frequencies in the dataset.
-Figure 3. Index sizes of indices with different key numbers built by the three methods on DBLP workload.
+VGGraph($max_{n}=3$, $c=0.5$)’s precision increase is insignificant.
 As the number of n-grams allowed in the index increases, starting from 2000, more indices from different configurations of FREE fall into this range.
-These configurations have low selectivity thresholds ($c<0.15$) and $n=2$ compared to the best index generated by FREE when $K=1000$, selecting n-grams that do not cover enough strings.
-This is demonstrated by the smaller sizes of these indices in Figure 3 even as number of keys increase, representing their short posting lists.
 The best configurations of FREE($\max_{n}=2$,$c=0.2$) achieve higher precision than BEST.
-The precision increases from 0.321 at 2000 keys to 0.535 at 5000 keys upper limit by FREE($\max_{n}=4$,$c=0.1$), without a significant increase in construction overhead.
-As expected, the index size grows with the number of keys.
-However, the improvement in workload performance is less significant compared to the change in precision.
-In the DBLP workload, each string in the dataset is short, resulting in very quick regex evaluation.
-Despite its precision, BEST achieves the best workload query performance.
-The index built by BEST has the smallest number of index keys, totaling 363.
-In comparison, a similar precision index by FREE requires 1440 keys.
-A larger index incurs greater index lookup overhead, which is significant when the dataset consists of short strings that match quickly after passing the filter.
-Another source of overhead for FREE comes from traversing its query plan trees, which are optimized for regexes with alternation (logical OR).
+The precision increases from 0.321 at 2000 keys to 0.524 at 4000 keys upper limit by FREE($\max_{n}=4$,$c=0.15$).
+Precision of VGGraph increases significantly from 0.167 to 0.725 as number of keys increases from 1014 to 3681.
+DBLP runs much faster with all index configurations than the baseline (i.e.
+using no indexes at all) of 99 seconds.
+Even when Trigrams indexes are built using only the first few trigrams—yielding low precision between 0.006 and 0.017, they still achieves performance improvement of $1.8\times$ to $3.6\times$.
+Looking at Figure 2(b), however, the improvement in workload performance is significant compared to the change in precision.
+The precision of FREE increases around 10 times as $K$ increases from 150 to 3000; its query time decreases by 41%.
+When $K=150$, the precision of BEST is significantly larger than FREE and LPMS in Figure 2(a), the performance difference for BEST and LPMS is not significantly different.
+In the DBLP workload, each string in the dataset is short, resulting in very quick regex evaluation, and therefore the query time reduction brought by precision increase is less significant.
+Trigrams’s precision increases as number of keys increases.
+When indexing all possible 28,275 trigrams, we can get a high precision of 0.97.
+However, the query time is 14.7s, similar to BEST with less than 500 keys.
+This suggests that indexing all trigrams is not cost-efficient in terms of storage.
 {customframe}
@@
 Generally, if the workload has a large number of queries that are not skewed (i.e., all very similar and covering a small subset of the dataset) and the strings in the dataset are not long, it is best to choose FREE.
+VGGraph significantly improve its precision only when the number of keys increase.
 When the query set is large, BEST and LPMS require significant computation time and memory, owing to their higher time and space complexities.
 Additionally, for a large-sized and balanced query set, frequent n-grams in the datasets are likely to be covered in the queries as well.
-Therefore, FREE benefits from these two aspects and outperforms the other two methods.
+Table 3. Index cost and query performance on Webpages.
+$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ GB $S_{I}$ MB PrecPrec
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 6.1.2. Workload Webpages
-Next, we examine a workload that is drastically different from the DBLP workload.
-As shown in Section 5.1, Webpages has a significantly smaller number of queries in its workload compared to DBLP.
-Each string in the Webpages dataset is an HTML file for a webpage, making the average string length the longest among all the workloads.
-We run the workloads using the three methods and summarize the experimental results in Section 6.1.2.
-We observed that the computational overhead of BEST is significantly reduced compared to DBLP due to the smaller query set.
-BEST achieves reasonable precision for performance improvement with a very small index size.
-However, since the query set is not representative, this small index may not benefit on unseen queries.
-Looking at the first row of Section 6.1.2, by selecting only 3 to 4 n-grams for indexing with BEST($c=0.1$), the index achieves a precision of 0.138.
-By manually early stopping FREE($\max_{n}=2$,$c=0.02$) at 5 n-grams selected, it achieves comparable precision of 0.123, while incurring around 20$\times$ lower time overhead.
-Table 4. Index cost and query performance on Webpages.
-$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ GB $S_{I}$ MB $Prec$
-5 BEST 422 8587 150 0.1 0.138
-FREE 532 428 148 0.1 0.123
-$\mathbf{1.7\times 10^{5}}$ BEST 422 8587 150 0.1 0.138
-FREE 353 1184 159 357.9 0.163
-$\mathbf{3.6\times 10^{5}}$ BEST 422 8587 150 0.1 0.138
-FREE 240 5965 240 2007.4 0.301
-$\mathbf{1.5\times 10^{6}}$ BEST 422 8587 150 0.1 0.138
-FREE 239 7480 233 524.9 0.302
-To attain a similar precision without early stopping, FREE($\max_{n}=2$,$c=0.5$) requires an index with more than 10000 times the number of keys and more than 3000 times the index size at $K=1.7\times 10^{5}$ The difference in the number of keys results from the different sources of n-grams considered during selection for the two methods.
-BEST chooses index keys from the intersection of n-grams in both the query set and the dataset, and a smaller number of queries means a smaller number of potential n-grams, effectively constraining the index size and computation overhead.
-On the other hand, FREE derives its index keys from the dataset, which contains many more distinct n-grams, leading to the selection of a large set of n-grams that meet the requirements defined by the specific configuration.
-Looking at the precision, we see that FREE, with a 0.025 higher precision than BEST, achieves a 1.2×1.2$\times$1.2 × overall query time speedup, reducing the time from 421.6 to 353.0 seconds.
+We evaluate DBLP, a workload with significantly fewer queries than DBLP, using HTML files from Webpages, which has the longest average string length among all workloads.
+Results are summarized in Section 6.1.1, with a baseline (no index) query time of 437s.
+At $K=5$, BEST($c=0.1$) selects just four n-grams and achieves 0.138 precision.
+A manually early-stopped FREE($\max_{n}=2$, $c=0.02$) with five n-grams yields a similar 0.123 precision.
+To match that precision without early stopping, FREE($\max_{n}=2$, $c=0.5$) requires $1.7\times 10^{5}$ keys.
+This difference stems from selection strategies: BEST chooses n-grams from the query–dataset intersection, limiting candidates when queries are few.
+In contrast, FREE selects from the dataset alone, resulting in a much larger candidate pool.
+Despite requiring more keys, FREE achieves 0.025 higher precision than BEST and delivers a 1.2× speedup in query time—reducing it from 421.6s to 353.0s as shown in Section 6.1.1.
 When the number of keys allowed $K$ increases from $1.7\times 10^{5}$ to $3.6\times 10^{5}$ , the precision of the index by FREE($\max_{n}=2$,$c=0.7$) increases significantly from 0.163 to 0.301.
-The index construction time also increases significantly.
-The query time of the best index built by FREE for this $K=3.6\times 10^{6}$ is 1.47×1.47$\times$1.47 × faster and 1.75×1.75$\times$1.75 × faster than the index built with the three n-grams selected by BEST.
-Compared to the DBLP workload, the precision increase in Webpages corresponds to a more significant workload runtime decrease from 353 seconds to 240 seconds.
-In this workload, each webpage is long, and the cost of regex evaluation on one string is high.
-Therefore, eliminating more strings with the index significantly improves the performance.
-These advantages might not justify the large index size compared to BEST for this specific workload.
-However, it does make the index more robust to unseen queries on the same dataset.
-LPMS did not finish for this dataset within the set time frame due to the large average number of characters per line of the dataset.
-Unlike BEST where the data structures for intermediate calculation are adjacency lists due to the sparseness Assumption Assumption BEST-4, LPMS builds matrices of size $|Q|\times|G|$ as input to integer program solver, where the actual computational and space overhead is too large.
+At $K=3.6\times 10^{5}$, FREE delivers a query time that is 1.47× faster than its own earlier version and 1.75× faster than BEST.
+Compared to the overall precision and query time trend for DBLP workload in Figure 2, the precision increase of FREE in Webpages corresponds to a more significant workload runtime decrease from 353 seconds to 240 seconds.
+Since each string in Webpages is a long HTML page, filtering more strings upfront yields greater performance benefits.
+While FREE’s larger index may not always justify its size relative to BEST, it offers better robustness to unseen queries on the same dataset.
+Due to lack of space, we defer the reader to Appendix B of the full paper for detailed experiments.
+VGGraph and LPMS did not finish for this dataset within the set time frame due to the large average length per string of the dataset.
 {customframe}
@@
 Two distinguishing characteristics of Prosite are: 1) it has the smallest alphabet size of 22, and 2) it has very short literal components in its the regex queries.
-Table 5. Index cost and query performance on Prosite.
-$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
-50 BEST 141.5 400 0.931 9.4 0.00826
-FREE 154.0 2 0.227 1.5 0.00651
-LPMS 151.8 14 0.263 26.2 0.00708
-100 BEST 139.7 856 0.967 14.9 0.0089
-FREE 153.6 2 0.242 5.7 0.00652
-LPMS 151.8 14 0.263 26.2 0.00708
-500 BEST 139.7 856 0.967 14.9 0.00890
-FREE 150.8 3 0.595 141.4 0.00687
-LPMS 151.8 14 0.263 26.2 0.00708
-Section 6.1.3 shows the results.
-BEST performs the best for this workload, generating index with highest precision while using a small number of keys of 364.
+The short literal components gives little opportunity for n-gram index improvement, and all methods created indices with low precision and performance improvement.
+In fact, the baseline workload runtime is 154 seconds.
+(a) Precision
+(b) Query Time $T_{Q}$ (s)
+Figure 3. Prosite.
+BEST is the most performant for this workload, when $K$ is small, generating index with highest precision while using only 364 keys.
 The small alphabet size and small literal size per query make the number of possible n-grams considered in each calculation iteration small, and thus drastically reduce the index construction runtime overhead, which is the biggest advantage of BEST in other workloads.
+Similarly, VGGraph which adapts a similar cover strategy, achieves the best precision as $K$ gets larger.
 When indexing with 50 keys, FREE($\max_{n}=2$,$c=0.15$) generates an index with a higher precision of 0.00651, while taking the least computational overhead and smallest index size.
 LPMS-D, taking slightly longer to select the n-grams and generate the index, achieves a precision of 0.00708 within 14 seconds.
-It achieves similar query performance compared to the index by FREE, although the index size is significantly larger, as it selects more frequent n-grams as keys.
 BEST($c=0.7$) achieves the highest precision, 0.00826, among the three methods.
 Its query time is also more than 10 seconds lower than the other two methods.
-However, it requires 29×29$\times$29 × the index construction time compared to LPMS and 200×200$\times$200 × compared to FREE.
-It also consumes significantly more compute memory, nearly 3×3$\times$3 × higher than that of FREE and LPMS when $K=50$ and $K=100$.
 Overall index precisions are low for Prosite workload, as the average length of literals is very short in the regexes.
-When $K=100$, the index by FREE($\max_{n}=2$,$c=0.2$) has almost identical precision with the 50-key index constructed by the same method.
-The workload running time for LPMS-D is slightly faster than FREE, with a similar compute memory.
+When $K=100$, FREE, LPMS, and Trigrams does not lead to much improvement.
 Index by BEST($c=0.7$) has a higher precision of 0.0089, resulting in a more than 10 seconds workload performance improvement over other methods.
-BEST takes 856 seconds to construct the index, much longer than other methods, but having smaller index size than LPMS.
-As the key size constraint loosens to $K=500$, FREE selects an index which only marginally decreases its overall workload running time from 153.6 seconds to 150.8 seconds, while using 395 keys (330 more keys than when $K=100$).
-The size of the index increases to 24.6×24.6$\times$24.6 × compared to the index built by FREE when $K=100$, and is much larger than the index sizes of other methods.
-However, its precision is still less than that of BEST and LPMS with a much smaller number of keys.
-The results align with the characteristics of the methods: BEST trades off n-gram selection time for higher precision n-grams tailored to the workload, while LPMS approximates the results of BEST by trading off some precision for lower index construction overhead.
-When comparing the index construction memory and time overhead to the results for the same $K$ in DBLP, we observe that for Prosite, BEST has significantly smaller overhead that is not proportional to the dataset size difference.
-In Prosite, the possible literal set is small due to the limited alphabet size and short query literal sizes.
-Consequently, the computation time and space are significantly reduced in each step when BEST needs to count the number of lines for each n-gram.
+VGGraph ($max_{n}=6$, $c=0.7$) also increases slightly to 0.0732.
+The precision increments for all methods does not corresponds to significant performance improvement as shown in the bars in Figure 3(b).
+As the key size constraint relaxes to $K=500$, selected n-grams of VGGraph ($max_{n}=2$, $c=0.7$) achieves the highest precision of 0.0987, and the significant precision improvement from the previous $K$ led to a query time reduction to 128.5 seconds, $1.13\times$ query time reduction from the baseline.
+Since the literal lengths are small in this workload, the effectiveness of each n-gram is much more important for effective filtering.
+Therefore, n-gram sets selected by dataset-based methods, FREE and Trigrams, are not effective in filtering.
 {customframe}
@@
 For workloads where query literals are short and/or the alphabet size is small, BEST can select an n-gram set with high filtering precision while incurring reasonable computational and storage overhead.
-This advantage arises from the decreased number of potential n-grams to consider.
+This advantage arises from the smaller n-gram candidate set.
+VGGraph outperform other methods in terms of precision as number of n-grams scales up.
 FREE does not benefit from this advantage since it only looks at n-grams in the dataset.
 LPMS benefits less from the characteristics of the workload due to the fixed overhead for constructing the integer program.
+(a) Precision
+(b) Query Time $T_{Q}$ (s)
+Figure 4. US-Acc.
 6.1.4. Workload US-Acc
@@
 Each string in the US-Acc dataset is much shorter.
-Table 6. Index cost and query performance on US-Acc.
-$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
-30 BEST 1.451 63 2.204 84.80 0.510
-FREE 4.940 29 0.969 0.02 0.032
-LPMS 1.353 235 2.876 60.85 0.572
-1000 BEST 1.451 63 2.204 84.80 0.510
-FREE 1.708 31 3.235 1003.99 0.460
-LPMS 1.353 235 2.876 60.85 0.572
-5000 BEST 1.451 63 2.204 84.80 0.510
-FREE 1.297 46 3.199 1070.97 0.811
-LPMS 1.353 235 2.876 60.85 0.572
-Section 6.1.4 shows the results.
+Figure 4 shows the results.
+The methods that uses both query set and dataset as n-gram sources generate better query performance than the baseline query time of 4.8 seconds.
 LPMS-D performs the best for the US-Acc workload, achieving a precision of 0.572 using only 12 keys.
-Although it takes 3.7×3.7$\times$3.7 × the index computation time compared to BEST($c=0.7$) and slightly more computation space overhead, LPMS uses 0.4×0.4$\times$0.4 × less index space to achieve this higher precision.
 When $K$ increases to 1000, FREE generates an index with 837 keys and a precision of 0.46, which is much lower than BEST and LPMS with fewer than 30 keys.
 This is because FREE selects n-grams based solely on their selectivity in the dataset, omitting those n-grams that are frequent in both the regex query and the dataset.
-In fact, the configuration of FREE that generates the highest precision with fewer than 1000 keys uses a selectivity threshold $c=0.7$, the largest across all of our experiments.
+When $K$ increases to 5000, VGGraph ($max_{n}=3$, $c=0.7$) generates a index with 4247 keys and achieves an query time of 1.51 seconds that is comparable to other faster methods.
 {customframe}
@@
 For workloads where the query literals are very short and common in the dataset, LPMS can quickly locate a small set of n-grams with high precision.
-Both BEST and FREE remove n-grams that are frequent in the dataset, as they both adhere to Assumption Assumption FREE-1 that does not hold for the dataset.
+VGGraph, BEST, and FREE remove n-grams that are frequent in the dataset, as they both adhere to Assumption Assumption FREE-1 that does not hold for the dataset.
 Table 4. Index cost and query performance on SQL-Srvr.
-$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
-50 FREE 3457 1342 76.7 6643 0.000945
-LPMS 1421 4569 147.5 4474 0.009855
-5000 FREE 1473 1483 74.0 6643 0.002698
-LPMS 1421 4569 147.5 4474 0.009855
+$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$GB $S_{I}$MB PrecPrec
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 6.1.5. Workload SQL-Srvr
-The SQL-Srvr workload has dataset characteristics similar to the US-Acc workload, as it is generated from several formatted strings—in this case, system log reporting strings.
-We summarize the results in Section 6.1.4.
-The size of both the query set and dataset are much larger than those in US-Acc, and therefore, BEST failed to finish within the set time-frame.
-LPMS performs better than FREE for similar reasons: it considers the query set without overlooking common n-grams.
-When we limit the number of n-grams to 50, LPMS-D selects an n-gram set with a precision of $9.855\times 10^{-3}$, which is 10×10$\times$10 × higher than the index generated by FREE($\max_{n}=2$,$c=0.7$) with the same number of keys.
-This is because FREE blindly selects selective n-grams without considering the queries.
-Given the variability in the values of the formatted system logs (e.g., VM IDs, cluster names), FREE may index short n-grams from them, resulting in an index that hardly benefits the actual query set.
-Consequently, the query time for LPMS is 2.43×2.43$\times$2.43 × faster.
-Although LPMS takes 3.4×3.4$\times$3.4 × longer and requires 1.9×1.9$\times$1.9 × more construction space and time overhead, the index built by LPMS uses 1.48×1.48$\times$1.48 × less space.
-Considering that LPMS and FREE use the same inverted index implementation and they have the same number of keys (50), the only difference in their index sizes comes from the posting lists lengths, which are the IDs of the data records.
-This suggests that LPMS selects n-grams that are more selective in the dataset.
-Combined with its higher precision and faster query time, LPMS selects n-grams that are both more selective in the dataset and more common in the query set, making them more beneficial.
-The results remain consistent until the upper limit of the number of n-grams grows to 5000.
-At this point, FREE($\max_{n}=2$,$c=0.03$) selects an n-gram set of size 4643, increasing the precision to $2.68\times 10^{-3}$.
-This is achieved with only a slight increase in index construction overhead and nearly identical index size compared to indexing with 50 n-grams selected by LPMS.
-However, the precision and query time are still worse than those of LPMS with 50 n-grams.
+The SQL-Srvr workload, like US-Acc, consists of structured strings—specifically, system log reports—but with a much larger dataset and query set.
+BEST was unable to complete within the time limit.
+Results are summarized in Section 6.1.4.
+[…]
 {customframe}
@@
 This is due to LPMS’s design choice of not discarding any n-grams based on selectivity threshold.
-Table 8. Index cost and query performance on Synthetic.
-$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$ $S_{I}$ $Prec$
-20 BEST 0.829 3.398 0.042 0.091 0.2084
-FREE 0.190 0.027 0.016 0.446 0.2672
-LPMS 1.211 3.720 0.053 0.010 0.1479
-100 BEST 0.444 31.659 0.043 0.452 0.4793
-FREE 0.190 0.027 0.016 0.446 0.2672
-LPMS 1.063 3.791 0.052 0.016 0.1757
-300 BEST 0.061 90.661 0.043 1.125 0.6451
-FREE 0.064 0.032 0.015 1.122 0.6453
-LPMS 1.063 3.791 0.052 0.016 0.1757
-6.1.6. Robustness Test
-For this test, we use the synthetic workload with different sets of regexes for index construction and query matching to test the robustness of each method for unseen queries.
-We summarize the result in Section 6.1.5.
-For $K=20$, FREE($\max_{n}=2$, $c=0.7$) builds an index with 16 keys that achieves the highest precision of 0.267.
-It also has the lowest index construction overhead.
-BEST($c=0.2$) achieves slightly lower precision of 0.208, but the query time is $\times4$ more than that of FREE.
-LPMS-D performs the worst with the lowest precision and highest query time.
-Note that BEST and LPMS also have significantly higher computational overhead compared to FREE.
-As $K$ increases to 100, BEST($c=0.5$) achieves the highest precision.
-It’s index construction time also increases by $\approx$3 times to 31.7s.
-This is more than 1000×1000$\times$1000 × longer than FREE.
-Note that the query times for BEST are not the lowest among the three methods, although it has the highest precision.
-This is due to the non-uniformness in the data record lengths, which breaks Assumption Assumption BEST-2.
-As we increase to $K=300$, FREE($max_{n}=2$, $c=0.12$) achieves the highest precision of 0.6453.
-BEST($c=0.2$) achieves similar precision and similar overall query time compared to FREE, but with a much larger index construction time which is 2833×2833$\times$2833 × slower than FREE.
-In all cases, LPMS performs unfavorable to the other two methods, since the n-gram selection strategy of LPMS is solely based on the benefit calculation using the query set and the dataset, without any heuristic like the pruning step of BEST.
+6.1.6. Workload Enron
+The Enron workload contains long data strings in the dataset like Webpages, yet it has relatively small dataset size and larger query size.
+The lengths of literal components among queries are generally short.
+Although each data string is long, the small dataset size and the short literal components make the computation of majority of the methods within our set time limit.
+We run BEST, FREE, LPMS, VGGraph, and Trigrams on this dataset and summarize the results in Section 6.1.6.
+Without indexing, Enron runs in 126 seconds as a baseline.
+BEST ($c=0.7$) achieves highest precision for $K=10$.
+With a precision of 0.148, around $1.4\times$ higher than other methods, BEST uses $1.2\times$ less time for the workload querying.
+As $K$ increases to 100, precision of BEST increases by $2.2\times$ to 0.321, and the query time decreases by $1.5\times$.
+LPMS-D, having similar strategy as BEST, also achieves precision incremental as $K$ increases from 10 to 100 with significantly reduced precision.
+Enron has one of the highest regex match rate, where without filtering, the true positive rate per regex is already around 0.105, suggesting that VGGraph and Trigrams does not perform effective filtering.
+When number of keys selected for the index continue to increase, the query performance and precision fails to increase proportionally.
+This is due to the fact that literal components in the queries are fairly common in the data strings, getting filtered out due to the selectivity threshold earlier on.
 {customframe}
 Insights
-For scenarios when a unseen queries on a fixed dataset are expected, the strategy that takes advantage of the query set (e.g.
-BEST and LPMS) may fall short.
-FREE, selecting n-grams based on the dataset, is robust in this setting.
+BEST demonstrates exceptional performance in Enron, efficiently leveraging its optimization-based approach with smaller query sets and longer literals.
+VGGraph shows substantial performance enhancement at moderate index sizes.
+LPMS offers balanced precision and performance, while FREE and Trigrams indexing lag behind due to their less targeted selection methods.
+Table 5. Index cost and query performance on Enron.
+$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$GB $S_{I}$MB PrecPrec
+[…]
+[…]
+LPMS 124 79 3 16.3 0.107
+[…]
+[…]
+FREE 123 35 3 5.1 0.109
+LPMS 101 2104 13 256.2 0.149
+[…]
+[…]
+6.1.7. Workload ADX
+The ADX workload has the largest dataset and a very small query set.
+Trigrams method would not generate useful index if tight constraint is posed the number of n-grams selected, and will generate index too large to fit into memory if no constraint imposed.
+VGGraph did not finish for all configuration on this workload.
+This workload runs in 1663 seconds without indexing.
+We run FREE, BEST, and LPMS on ADX workload and summarize the result in Section 6.1.7.
+Due to the small query set size, and large and diverse data strings in the dataset, methods considering both query and dataset, BEST and LPMS achieves better precision and performance improvement for all key constraints $K$.
+Overall, LPMS-D achieves best precision of 0.328 and query performance of 221 seconds, $7.52\times$ faster than the baseline.
+BEST ($c=0.1$) also achieves $5.3\times$ to $6.5\times$ performance improvement with 10 to 25 keys.
+{customframe}
+Insights
+For ADX, a workload with simple patterns and very large dataset size, LPMS stands out by effectively balancing precision and computational overhead.
+BEST becomes computationally prohibitive due to its complexity, and Trigrams indexing is impractical at this scale.
+Table 6. Index cost and query performance on ADX.
+$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$GB $S_{I}$MB PrecPrec
+[…]
+[…]
+LPMS 373 2184 104 2799.0 0.113
+25 BEST 254 28912 219 650.6 0.216
+[…]
+[…]
+[…]
+[…]
+[…]
+6.2. Index Construction Overhead
+In this section, we aim to answer question Q.2 with a detailed analysis of the index construction overhead for FREE, BEST, LPMS, VGGraph, and Trigrams, based on their time and space usage observed during index construction.
+FREE demonstrates a significantly lower index construction time across various workloads due to its straightforward selectivity-based strategy and iterative prefix-free minimal n-gram selection approach.
+Specifically, for the Webpages workload, as shown in Section 6.1.1, FREE constructed an index with $1.7\times 10^{5}$ keys in merely 1184 seconds compared to BEST’s significantly longer 8587 seconds.
+FREE’s runtime space overhead remains relatively constant and low across workloads, generally proportional to the dataset size due to the prefix-free minimal set construction method.
+BEST has a high index construction overhead because it performs an exhaustive search to approximate optimal n-gram set through a utility-based set-cover strategy.
+Specifically, in the DBLP workload, BEST required 8762 seconds to construct an index of 363 keys, dramatically more than the sub-second construction time achieved by FREE for a comparable number of keys.
+The high overhead results from the auxiliary data structures mentioned in Section 4.2.2, required to maintain for time-efficient computation.
+In workloads with larger query sets or longer literals, such as Webpages in Section 6.1.1, BEST’s computation overhead increases and potentially exceeding practical time limits.
+LPMS has moderate to high index construction times, often higher than FREE but lower than BEST.
+For instance, in the SQL-Srvr workload results in Section 6.1.4, LPMS required 4569 seconds for constructing an index with 50 keys, significantly exceeding FREE’s 1342 seconds for an similar number of keys but less than BEST.
+LPMS’s overhead results from its formulation and solving of the linear program.
+VGGraph has an index construction overhead generally lower than BEST but higher than FREE.
+The overhead is attributed to its iterative evaluation of cost-efficiency ratios.
+For the SQL-Srvr workload in Section 6.1.4, VGGraph constructed indices in 45 seconds, considerably lower than BEST and LPMS that runs more than 2000 seconds but slightly higher than FREE (30s).
+Its overhead scales as the dataset and query complexity increase.
+Trigrams has minimal index construction overhead when limited to a small number of keys, but this overhead grows rapidly as more trigrams are indexed—especially without key constraints.
+Its cost is primarily driven by dataset size rather than query complexity or optimization.
+(a) Precision
+(b) Index Time $T_{I}$ (in seconds)
+Figure 5. Synthetic workload with $|D|=50,000$ and average query selectivity 0.02.
+6.3. Case Study: Workload Characteristics
+In this section, we perform an ablation study with Synthetic workload to evaluate how different regex indexing methods behave under varying query set sizes and query selectivity.
+For each configuration, we constrain the maximum number of n-grams ($K$) to 5000 and measured both precision and query runtime ($T_{Q}$).
+BEST failed to finish within the time limit.
+However, in the experimental result, varying regex selectivity does not result in a common trend on the indices built with different methods in workload performance or index construction.
+We will focus on the analysis of varying query size with fixed dataset size and regex query selectivity.
+Figure 5(a) illustrates that as the query set size increases from 1000 to 5000, FREE consistently maintained high precision.
+This can be attributed to the small alphabet size of the workload.
+LPMS and VGGraph also showed stable but lower precisions, with LPMS ranging around 0.95 and VGGraph at approximately 0.96.
+All three methods achieves similarly high precision.
+The precision of all methods slightly improve as query set size increased, as a larger query set provides more opportunities for n-gram filtering.
+Regarding index construction time ($T_{I}$), in Figure 5(b), FREE and VGGraph exhibited significantly lower overhead, maintaining fast index build times which is under 1 second across all query set sizes.
+LPMS required considerably more time as the query set size increased.
+LPMS is sensitive to query set growth, requires over 90 seconds at larger query sets.
+This matches our observation from Section 6.2, where LPMS and BEST incurred much larger index construction overhead.
 7. Learnings and Future Work
@@
 We summarize the insights in the decision flow chart in Figure 6 as a guideline for practitioners.
-This guideline recommends 1) choosing FREE for large and diverse workloads or when unseen queries are expected, 2) choosing BEST when precision is crucial for repeated queries and the number of candidate n-grams is small, and 3) choosing LPMS for formatted datasets or when query literals are common in the dataset.
+This guideline recommends 1) choosing FREE for large and diverse workloads or when unseen queries are expected, 2) choosing BEST when precision is crucial for repeated queries and the number of candidate n-grams is small, 3) choosing LPMS for formatted datasets or when query literals are common in the dataset, and 4) choosing VGGraph when precision requirement is high with low index space constraint.
 Based on this study, we also propose the following potential directions for future work in this area.
@@
 We also highlight the important open problems for regex indexing to spur further research in this area.
+Appendix A FREE Regex Query Plan Compiler
+In order to also handle regex queries literals in alternative strings, FREE include a simple regex query plan compiler that compiles the overall index lookup to a plan tree with only logical AND and OR operators.
+Specifically, for a given regex, the compiler identifies literals on the primary level and literals, and then build a tree with the literals.
+FREE then checks if each of the literals or parts of the literals are indexed in the index, removing subtrees with no literals indexed, and substituting literal nodes with index keys, which are n-grams in the index.
+We provide an example index search query plan in Figure 7 for the example regex, <ahref=("|’).*ZZZ\.
+pdf("|’)>, assuming that only ", , Z, and pdf are indexed.
+The literal tree shown in Figure 7(a) is built solely by extracting all literals from the regex.
+Referring to the index, we notice that <ahref= and > are not indexed, and two n-grams in the literal ZZZ.pdf, Z and pdf are indexed.
+We simplify the plan to get Figure 7(b).
+The evaluation of the index search plan tree then is a depth-first evaluation, where AND is a set intersection operation of two posting lists and OR is a set union operation of two posting lists.
+Appendix B Robustness Test
+For this test, we use the synthetic workload with different sets of regexes for index construction and query matching to test the robustness of each method for unseen queries.
+To test the robustness of the n-gram selection methods in handling unseen queries, we build off the synthetic dataset in the original LPMS (Tsang and Chawla, 2011) paper.
+The dataset consists of 5000 strings each constructed with alphabet ‘A-P’.
+The string size follows a geometric distribution with $p=1/32$.
+The index building query set is generated by randomly selecting 10% from the dataset, and a random slice is used to create a regex lit1regexlit2.
+lit1 has 1-5 characters, lit2 has 0-5 characters, and regex matches any $m$ characters where $1\leq m\leq 50$.
+During the query phase, we generate regexes of a similar format from 2% of the data records.
+We summarize the result in Appendix B.
+Overall, FREE demonstrates robustness in this experiment, achieving highest precision with the lowest computational overhead.
+This can be attributed to the fact that FREE selects n-grams independent of the query set, and thus adapt smoothly to unseen regexes in query time.
+For $K=20$, FREE($\max_{n}=2$, $c=0.7$) builds an index with 16 keys that achieves the highest precision of 0.267.
+It also has the lowest index construction overhead.
+BEST($c=0.2$) achieves slightly lower precision of 0.208, but the query time is $4.4\times$ more than that of FREE.
+LPMS-D performs the worst with the lowest precision and highest query time.
+Note that BEST and LPMS also have significantly higher computational overhead compared to FREE.
+As $K$ increases to 100, BEST($c=0.5$) achieves the highest precision.
+It’s index construction time also increases by $\approx$3 times to 31.7s.
+This is more than $1000\times$ longer than FREE.
+Note that the query times for BEST are not the lowest among the three methods, although it has the highest precision.
+This is due to the non-uniformness in the data record lengths, which breaks Assumption Assumption BEST-2.
+As we increase to $K=300$, FREE($max_{n}=2$, $c=0.12$) achieves the highest precision of 0.6453.
+BEST($c=0.2$) achieves similar precision and similar overall query time compared to FREE, but with a much larger index construction time which is $2833\times$ slower than FREE.
+In all cases, LPMS performs unfavorable to the other two methods, since the n-gram selection strategy of LPMS is solely based on the benefit calculation using the query set and the dataset, without any heuristic like the pruning step of BEST.
+{customframe}
+Insights
+For scenarios when a unseen queries on a fixed dataset are expected, the strategy that takes advantage of the query set (e.g.
+BEST and LPMS) may fall short.
+FREE, selecting n-grams based on the dataset, is robust in this setting.
+Table 7. Index cost and query performance on Robust.
+$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$GB $S_{I}$MB PrecPrec
+20 BEST 0.829 3.398 0.042 0.091 0.2084
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]